### v1

# Quantum-enabled optical large-baseline interferometry: applications, protocols and feasibility

Abstract
Optical Very Long Baseline Interferometry offers the potential for unprecedented angular resolution in both astronomical imaging and precision measurements. Classical approaches, however, face significant limitations due to photon loss, background noise, and the requirements for dynamical delay lines over large distances. This document surveys recent developments in quantum-enabled VLBI, which aim to address these challenges using entanglement-assisted protocols, quantum memory storage, and nonlocal measurement techniques. While its application to astronomy is well known, we also examine how these techniques may be extended to geodesy – specifically, the monitoring of Earth’s rotation. Particular attention is given to quantum-enhanced telescope architectures, including repeater-based long-baseline interferometry and quantum error-corrected encoding schemes, which offer a pathway toward high-fidelity optical VLBI. To aid the discussion, we also compare specifications for key enabling technologies to current state-of-the-art experimental components, including switching rates, gate times, entanglement distribution rates, and memory lifetimes. By integrating quantum technologies, future interferometric networks may achieve diffraction-limited imaging at optical and near-infrared wavelengths, surpassing the constraints of classical techniques and enabling new precision tests of astrophysical and fundamental physics phenomena.
I Introduction
The performance of an imaging system or interferometer is limited by diffraction: the resolution is proportional to its aperture and inversely proportional to the wavelength $\lambda$ of the light. Together, these place a fundamental limit on how well one can image the objects of interest. This is especially true for light originating in astronomical objects, where we cannot illuminate the objects of interest, and we are restricted to analysing the light that reaches us.
Several challenges hinder the progress in building large-baseline optical interferometers, including the weight of the instrument itself, photon loss, background noise, and the requirements for long dynamical delay lines. These factors ultimately limit the achievable distance between telescope sites. Quantum technologies can help bypass transmission losses: using quantum memories and entanglement, we can replace the direct optical links, allowing for much larger distances.
Among the key scientific applications of optical Very Large Baseline Interferometry (VLBI), two stand out. First, VLBI offers a path toward high angular resolution astronomical imaging, particularly in regimes that are inaccessible to conventional telescopes. Second, and less commonly considered, optical VLBI can achieve unprecedented spatial and temporal resolution in geodetic metrology, such as monitoring Earth’s rotation, tectonic motion, and lunar libration.
Today’s most advanced interferometric imaging systems operate at radio and microwave wavelengths. Instruments such as the Event Horizon Telescope (EHT) achieve resolutions of tens of microarcseconds by combining a maximum baseline of $\sim$12,000 km (roughly the Earth’s diameter) with a typical wavelength of 3 mm. Shifting to optical wavelengths can, in principle, increase resolution by several orders of magnitude. For instance, an optical system operating at $\sim$600 nm would outperform the EHT in resolution with a baseline as short as 2.4 km. At 1550 nm (a standard C-band telecom wavelength), the equivalent threshold is 6.2 km.
II Astronomical imaging
2019 saw a dramatic demonstration of the image resolution as limited by the size of the aperture of the imaging system. The Event Horizon Telescope collaboration combined telescopes from all over the world into a single imaging system with an aperture roughly the size of the Earth. This collective effort resulted in the first-ever image of a black hole and its surroundings.
The underlying physics of the black hole imaging setup is purely classical: the array operates at radio frequencies, where both the amplitude and phase of the received electromagnetic field can be directly measured. The data is then post-processed to reconstruct an image. Moving to optical frequencies, however, introduces significant complications: even the fastest electronics cannot directly measure the oscillations of the electric field at optical frequencies. Nevertheless, if these complications can be overcome, the potential gains in resolution are substantial. Optical VLBI offers a 3-5 orders of magnitude improvement in resolution over radio-based systems, depending on the wavelength and baseline. This opens the door to imaging astrophysical structures that are beyond the reach of our current imaging systems.
Optical VLBI refers to interferometric regimes where real-time calibration of optical path lengths is not feasible, necessitating the use of closure phases and post-processed visibilities [7]. Instruments like CHARA [8] and VLTI [9] operate in this regime, and future concepts such as the Big Fringe Telescope [10] aim to extend it. Our work does not propose optical VLBI as a new idea, but rather explores how quantum-enhanced protocols might overcome key technical limitations, particularly in photon loss and scalability.
The advantage of a quantum approach is not merely improved angular resolution based on extending the baseline, but achieving the best precision in complex visibility measurements. Many science cases—such as detecting exoplanets, imaging stellar surfaces, or tracking relativistic orbital precession—rely on high signal-to-noise measurements rather than resolving power alone.
Current optical interferometers face severe constraints on scalability and baseline length, due to the need for physical delay lines and lossy multi-way beam combination. Quantum architectures offer a route to nonlocal beam combination by connecting telescopes without direct optical interference. If deployed on the Moon, we can exploit the natural vacuum as a beam guide for long-baseline operation without atmospheric distortion.
These approaches are complementary to other classical strategies such as heterodyne interferometry [11], and intensity interferometry [12]. However, the quantum approach uniquely targets large-baseline, high-precision measurements and scalable architecture without the need for formation flying or complex delay-line infrastructure.
In the remainder of this section discuss some science cases for using large-baseline optical interferometry, including imaging of exoplanets, tracking the stars around black holes, and measuring separations of double-star systems.
II.0.1 Imaging exoplanets
Exoplanet Semi-Major Axis (AU) Distance (pc) Separation (arcseconds)
Proxima Centauri b 0.0485 1.301 0.037
Barnard’s Star b 0.0406 1.834 0.022
Ross 128 b 0.0496 3.374 0.015
Luyten’s Star b 0.0911 3.785 0.024
Wolf 1061 c 0.084 4.287 0.020
Table 1: Angular separation of nearby exoplanets from their parent stars.
Exoplanets present a formidable observational challenge: they are both extremely small in angular size and orders of magnitude fainter than their host stars. Direct imaging of these objects requires two key capabilities:
  • High angular resolution, required to spatially separate the planet from the glare of the star.
  • High visibility (contrast), required to detect the planet’s faint signal against the much brighter stellar companion.
Optical VLBI offers a potential solution to both challenges. At a wavelength of 600 nm and a baseline of 100 km, the achievable angular resolution is approximately 1.24 microarcseconds ($\upmu$as)—well beyond the reach of existing instruments, including the Event Horizon Telescope. At this resolution, it becomes conceivable to directly image spatially resolved features in the atmospheres of nearby exoplanets, opening new possibilities for studying their composition, structure, and potentially even weather systems. We show the separation of the star from the exoplanet in Table 1.
II.0.2 Tracking stars orbiting black holes
Studying the orbits of stars in the strong gravitational fields near black holes provides a unique opportunity to test Einstein’s General Relativity (GR) in regimes that cannot be accessed in laboratory experiments. In Newtonian gravity, the orbit of a test particle around a massive body is a closed ellipse, as described by Kepler’s first law. However, GR predicts that the curvature of spacetime near massive objects causes the orbit to precess — a phenomenon known as Schwarzschild precession. The precession angle per orbit is given by
$\Delta\Phi=6\pi\frac{GM}{a(1-e^{2})c^{2}},$ (1)
where $G$ is the gravitational constant, $M$ is the mass of the central object, $a$ is the semi-major axis, $e$ is the eccentricity of the orbit, and $c$ is the speed of light. This effect, which is first order in $(v/c)^{2}\sim 10^{-4}$, has been experimentally confirmed [13].
In cases where the black hole is spinning, there is an additional relativistic effect known as frame dragging, or the Lense–Thirring effect: the rotation of the black hole causes spacetime itself to twist (Fig. 1), subtly altering the motion of nearby stars. The corresponding precession per orbit is given by
$\Delta\Phi_{\text{LT}}=2\chi\left(\frac{R_{S}}{a(1-e^{2})}\right)^{1.5},$ (2)
where $\chi$ is the dimensionless spin parameter of the black hole (ranging from 0 to 1), and $R_{S}$ is the Schwarzschild radius. For the star S2, which orbits the supermassive black hole at the centre of our galaxy, the Lense–Thirring precession is of the order of 0.05 arcminutes per orbit — about 240 times smaller than the Schwarzschild precession [13]. Resolving such small effects requires extremely high angular resolution. The GRAVITY collaboration (see, e.g., Refs. [14, 15, 16]) currently achieves near-infrared resolution of 3 milli-arcseconds and astrometric precision of 10–100 $\upmu$as, and a baseline of 130 m. To detect the Lense–Thirring precession at comparable wavelengths (near infrared), a baseline of at least $130\times 240=32$ km would be required.
Figure 1: The Lense–Thirring effect: a star’s orbit precesses due to the frame-dragging caused by a rotating massive object, such as a spinning black hole.
II.0.3 Measuring the stellar separation of A. Centauri.
Alpha Centauri consists of two Sun-like stars (Alpha Centauri A and B) in a close binary orbit. The angular separation of the two main stars in the Alpha Centauri binary system — Alpha Centauri A and Alpha Centauri B — varies over their 79.91-year orbit. It ranges from about 2 arcseconds at closest approach (periastron) to about 22 arcseconds at their farthest (apastron). By studying their motion, we can precisely measure their masses using Kepler’s laws and Newtonian mechanics. This helps in stellar modelling, as mass is a key determinant of a star’s lifetime and properties. Tracking deviations from expected orbital motion can reveal exoplanets, especially Earth-like planets in the habitable zone.
III Geodesy
Figure 2: Geometry for using VLBI for geodesy. The position of the reference star is assumed fixed, and the goal is to measure the baseline vector $\bm{B}$.
VLBI is a geometric technique; it measures the time difference between the arrival at two Earth-based antennas emitted by a distant quasar. The basic geometric delay $\tau$ measured by two radio telescopes with Earth-based positions $r_{1}$ and $r_{2}$ is given by
$\tau=-\frac{\bm{\mathrm{B}}\cdot s}{c}$ (3)
where $\bm{\mathrm{B}}$ is the baseline vector $\bm{\mathrm{B}}=r_{2}-r_{1}$, ${}s$ is the radio source vector, and $c$ is the speed of light, see Fig 2.
Using large numbers of time difference measurements from many quasars observed with a global network of antennas, VLBI determines the inertial reference frame defined by the quasars and simultaneously, the precise positions of the antennas. Since the antennas are fixed to the Earth, their locations track the instantaneous orientation of the Earth in the inertial reference frame. Relative changes in the antenna locations from a series of measurements indicate tectonic plate motion, regional deformation, and local uplift or subsidence.
The core idea is the precise timing of light signals:
  1. 1.
    When a quasar emits EM waves, they arrive at different telescopes at slightly different times due to Earth’s shape and rotation.
  2. 2.
    The exact time delay between telescopes depends on the Earth’s rotation state at that moment.
  3. 3.
    By comparing the signals and modelling their delay based on Earth’s motion, scientists can infer tiny variations in rotation speed and axis position.
Geodetic VLBI plays a crucial role in geodesy, enabling precise measurements of the Earth’s radio telescope positions and long-term motion of the tectonic plates. Currently, the tectonic motion is measured with accuracy up to 0.1 mm/year, though the seasonal variations of the individual radio telescopes’ positions are also monitored [17]. Some of the radio telescopes were affected by strong earthquakes (Japan, 2010; Chile, 2011) [18]. Hence, the co-seismic and post-seismic displacement can be monitored. A global catastrophe, like the Sumatra Boxing Day earthquake on 26-Dec-2004, may cause a global shift in positions of all geodetic sites around the Earth with an amplitude of $\sim$1 mm. In addition, geodetic VLBI studies the post-glacial uplift in Fenno-Scandinavia, the northern part of North America, and Antarctica. The global melting of glaciers causes a dramatic change in the immediate positions of radio telescopes, and changes the shape of the planet due to the secular shift of the inertia momentum.
III.1 Earth rotation
III.1.1 Time Scale
One of the primary applications of geodetic VLBI is the daily monitoring of UT1 (Universal Time 1), which tracks Earth’s rotational time. Earth’s rotation is not constant due to interactions with atmospheric winds, ocean currents, and the fluid outer core. These variations necessitate daily UT1 updates to account for fluctuations in the length of the day. The length of day (LOD) has increased by 37 sec since 1962 (Fig 3).
Figure 3: Increasing of the length of day (LOD) since 1962.
Another key application is long-term drift monitoring, which allows scientists to detect millisecond-scale changes in Earth’s rotation over decades. Such long-term observations reveal that Earth’s rotation is gradually slowing, with an average deceleration of 1.7 milliseconds per century due to tidal friction from the Moon’s gravitational influence.
The leap second problem arises from the need to synchronise two fundamentally different timekeeping systems: Atomic Time (TAI/UTC) and Astronomical Time (UT1). Atomic Time, based on hydrogen and caesium atomic clocks, is highly stable and defines the SI second with unparalleled precision. In contrast, Astronomical Time (UT1) is tied to Earth’s rotation, which varies unpredictably. As a result, UT1 drifts relative to atomic time, necessitating the periodic introduction of leap seconds to maintain synchronisation. In the 1990s, the leap second was introduced almost every year. Although the Earth rotation deceleration nowadays is not so high and the leap second is introduced rarely (the last time was in 2015), the situation may change again and needs to be monitored continually [19].
Leap seconds introduce significant challenges for computing, satellite navigation, and global timekeeping systems. Since leap seconds occur irregularly and unpredictably, they disrupt time-dependent operations, requiring software updates and manual interventions. To mitigate these issues, discussions are ongoing regarding the potential elimination or redefinition of leap seconds in the future [20]. Optical VLBI offers substantial improvement in the temporal resolution of UT1 measurements, enabling detection of changes in Earth’s rotation on shorter and more precise timescales.
III.1.2 Polar motion
Polar motion is the motion of the instantaneous Earth rotation axis around the geographic pole. The regular motion consists of two major components, namely the annual rotation with a period of approximately 365 days and the Chandler wobble with a period of approximately 430 days. The Chandler wobble is a free rotation effect with variable period and amplitude. Currently, it has disappeared from the total Earth’s pole motion (presumably temporarily) for unknown reasons. VLBI helps in tracking polar motion, as the Earth’s rotational axis drifts slightly due to mass redistributions, such as the melting of ice sheets and shifts in the planet’s mantle [21][22]. Optical VLBI can enhance our ability to track polar motion with finer spatial precision due to the shorter wavelengths involved, potentially improving sensitivity to small-scale mass redistributions.
III.1.3 Precession and nutation
The Earth’s axis draws a pattern in the sky. This pattern splits into two effects, precession and nutation. Precession has a period of 26,000 years and traces out a cone with an opening angle of $23.5^{\circ}$. The slow motion of the axis leads to a displacement of all apparent positions of the celestial objects of 50 arcsec every year. Nutation is a smaller signal with a major amplitude of about 9 arcsec and a period of 18.6 years, caused by the gravitational attraction of the Moon. Analysis of the numerous subtle effects in the nutation, e.g. free core nutation (FCN) and the free inner core nutation (FICN), helps to gain insights into the Earth’s deep interior dynamics and composition [23, 24, 25].
III.2 Astrometry
Geodetic VLBI produces the International Celestial Reference Frame (ICRF) by measuring the positions of several thousand strong radio sources. The current International Celestial Reference Frame (ICRF3) adopted by the International Astronomical Union is based on the positions of 4536 radio sources observed between 1979 and 2018, and 303 radio sources define the fixed position of the Earth’s fundamental rotation axis with an uncertainty of 30 $\upmu$as [26]. Detection errors of the individual positions of radio sources vary considerably due to an uneven number of observations, from 6 $\upmu$as to a few mas. Recent observations after 2018 increased the number of monitored radio sources to approximately 5600 [27].
The main ICRF3 reference catalogue is based on observations at a frequency of 8.4 GHz (the X band). In parallel, there are additional catalogues for frequencies of 24 GHz and 32 GHz with fewer objects. These catalogues are important for tracking deep space missions [28]. If optical VLBI is used, the enhancement is a finer determination of source positions in the celestial reference frame, with potential uncertainties reduced by 3-5 orders of magnitude compared to radio VLBI.
IV The model and parameter estimation
High-resolution astronomical imaging and high-precision geodesy are fundamentally linked by their reliance on the same core technology: interferometry. In astronomical imaging, interferometers are used to resolve fine spatial structure by coherently combining light collected from a distant source. In geodesy, the principle is inverted: rather than using interferometry to study a celestial target, the celestial source serves as a fixed reference, and the interferometric measurement is used to track the motion of the Earth itself. In both contexts, the required technology is essentially the same, therefore, in this work, we adopt a common physical model for the stellar source. To facilitate analytic comparison and benchmarking, we focus on the well-understood two-mode model, for which the quantum Fisher information is readily computable and serves as a natural metric of the resolution.
Consider a single frequency band whose two-mode continuous variable quantum description is given by a mean vector and covariance matrix of the form [29, 30, 5],
$\bm{\mathrm{r}}_{\star}$ $:=(\begin{array}[]{cccc}0&0&0&0\end{array})^{T},$ (5)
$\bm{\mathrm{\sigma}}_{\star}$ $:=\left(\begin{array}[]{cccc}\epsilon+1&0&\gamma\epsilon\cos\phi&-\gamma\epsilon\sin\phi\\ 0&\epsilon+1&\gamma\epsilon\sin\phi&\gamma\epsilon\cos\phi\\ \gamma\epsilon\cos\phi&\gamma\epsilon\sin\phi&\epsilon+1&0\\ -\gamma\epsilon\sin\phi&\gamma\epsilon\cos\phi&0&\epsilon+1\end{array}\right),$ (10)
where we have used the quadrature ordering $(q_{A},p_{A},q_{B},p_{B})$, with subscripts referring to Alice ($A$) and Bob ($B$).
In the limit that $\epsilon\ll 1$, which is typically true for astronomy, we can describe the optical state by the density matrix,
$\rho_{\star}\approx$ $(1-\epsilon)\ket{\text{vac},\text{vac}}\bra{\text{vac},\text{vac}}_{AB}+$
$\epsilon\left(\frac{1+\gamma}{2}\right)\ket{\psi_{+}^{\phi}}\bra{\psi_{+}^{\phi}}+\epsilon\left(\frac{1-\gamma}{2}\right)\ket{\psi_{-}^{\phi}}\bra{\psi_{-}^{\phi}}$ | (11)
[…]
If we were to divide a bandwidth spectrum (centred at 500 THz or 600 nm) of $\Delta\lambda=10$ nm into 100 MHz narrow bands, this is approximately 8300 bands; the photon will have a coherence time of approximately 10 ns with a coherence length 3 m; if instead we divide this into 1 GHz bands (830 bands), then the coherence time will be 1 ns, and the coherence length will be 30 cm. In the literature, the bandwidth of the stellar source is often quoted as either frequency or wavelength; we include Table 4 for reference and ease of converting between them.
For imaging, the parameters of interest are $\phi$ and $\gamma$, where $\phi\in[0,2\pi)$ is related to the location of the sources, and $\gamma\in[0,1]$ is proportional to the Fourier transform of the intensity distribution via the van Cittert-Zernike theorem [29].
The ultimate precision in parameter estimation is specified by the quantum Cramér–Rao bound [31, 32] (see also [33, 34]). For estimation of a parameter $\theta$ encoded into a quantum state $\hat{\rho}_{\theta}$, the Cramér–Rao bound sets a lower bound on the variance $(\Delta\theta)^{2}=\langle\theta^{2}\rangle-\langle\theta\rangle^{2}$ of any unbiased estimator $\theta$. For unbiased estimators, the quantum Cramér–Rao bound establishes that
$(\Delta\theta)^{2}\geqslant\frac{1}{NJ_{\theta}(\hat{\rho}_{\theta})}\,,$ (12)
where $N$ is the number of copies of $\hat{\rho}_{\theta}$ used and $J_{\theta}$ is the quantum Fisher information (QFI) associated with the state $\hat{\rho}_{\theta}$.
If there are multiple parameters we want to estimate, where $\bm{\mathrm{\theta}}=(\theta_{1},\theta_{2},\dots)$, we can define a QFI matrix $\bm{\mathrm{J}}$ that quantifies not only the QFI for each parameter (diagonal components) but also for correlated parameters (off-diagonal components). The matrix elements are given by
$J_{jk}:=\frac{1}{2}\text{Tr}[\hat{\rho}_{\bm{\mathrm{\theta}}}(\hat{L}_{j}\hat{L}_{k}+\hat{L}_{k}\hat{L}_{j})],$ (13)
where $\hat{L}_{j}$ is the symmetric logarithmic derivative with respect to $\theta_{j}$ [35].
The inverse of the QFI matrix provides a lower bound on the covariance matrix $\big{[}\operatorname{Cov}(\bm{\mathrm{\theta}})\big{]}_{jk}=\braket{\theta_{j}\theta_{k}}-\braket{\theta_{j}}\braket{\theta_{k}}$,
$\operatorname{Cov}(\bm{\mathrm{\theta}})\geq\frac{1}{N}\bm{\mathrm{J}}^{-1}.$ (14)
The QFI matrix elements for the incoming stellar state are found to be
$J_{\phi}$ $=\frac{2\gamma^{2}\epsilon}{2+\epsilon(1-\gamma^{2})},$ (15a)
$J_{\gamma}$ $=\frac{2\epsilon\left(2+\epsilon+\epsilon\gamma^{2}\right)}{\left(1-\gamma^{2}\right)\left(4+4\epsilon+\epsilon^{2}\left(1-\gamma^{2}\right)\right)},$ (15b)
$J_{\phi\gamma}$ $=0$ (15c)
If we take the equivalent trace norm of the QFI matrix,
$||\mathbf{J}||=J_{\phi}+J_{\gamma}$
$=2\epsilon\left(\frac{\gamma^{2}\epsilon+\epsilon+2}{\left(\gamma^{2}-1\right)\left(\left(\gamma^{2}-1\right)\epsilon^{2}-4\epsilon-4\right)}+\frac{\gamma^{2}}{\gamma^{2}(-\epsilon)+\epsilon+2}\right)$
$=\frac{(1+\gamma^{2}-\gamma^{4})}{1-\gamma^{2}}\epsilon-\frac{1}{2}\left(1+\gamma^{2}-\gamma^{4}\right)\epsilon^{2}+O(\epsilon^{3})$ (16)
Note that the above expression can be significantly larger than $\epsilon$ for $\gamma\rightarrow 1$. It’s also perhaps unfair to lump them together because $J_{\phi}/\epsilon$ is capped by $1$, whereas $J_{\gamma}/\epsilon$ can, in principle, approach infinity.
For geodesy, the task is to measure the distance between two telescope sites, given that a stable astronomical source is at a known location (Fig. 2). The most accurate way of measuring this distance is by estimating the optical path length difference between the two sites, which manifests as a phase shift $\phi$.
      Requirements       Direct detection       quantum-enabled
      Phase reference and phase stabilisation
      Variable delay lines
      (fast?) time-bin multiplexing
      High cooperativity optical cavities
      High-fidelity light-matter coupling/gates
      Entanglement distribution/ distillation
      Long-lived quantum memories
      Multi-system quantum gates
Table 2: The technology required for direct interferometry vs quantum VLBI’s.
Now, since the source is a point source, the quantum state shared between the two stations as
|$\ket{\psi}$ $=\frac{1}{\sqrt{2}}(\ket{01}+e^{i\phi}\ket{10}),$
$\phi$ $=\omega\tau,$
$=\frac{B\sin\theta}{\lambda}.$ (17)
From quantum metrology, we know that the QFI of $\phi$ is equal to 1; i.e. $\Delta\phi\approx 1$. Here we assume $\theta$ to be fixed, and $\lambda$ is also fixed. This means
$\Delta B\geq\frac{\sin\theta}{\lambda}\frac{\Delta\phi}{\sqrt{n}},$ (18)
where $n$ is the number of photons received.
The relation in Eq. (18) reveals a key advantage of operating at optical wavelengths: since the uncertainty in the baseline measurement scales inversely with $\lambda$, moving from radio to optical can potentially yield a substantial gain in precision. For the same number of detected photons and observational geometry, shorter wavelengths directly translate into increased precision of geodetic measurements.
V Technological requirements of OVLBI
Realising quantum-enabled optical VLBI involves a multi-layered technological framework that integrates several key components (see Table 2):
  1. 1.
    The entire setup needs to be phase-stabilised to within a small fraction of the wavelength. A shared phase reference is needed.
  2. 2.
    Collect the starlight and time-multiplex them into different bins (see Fig. 4)
  3. 3.
    Each time bin is coupled to a cavity, the cavity is used for enhancing light-atom interaction.
  4. 4.
    The collected starlight is absorbed into quantum memories. Non-local measurement is performed between quantum memories of the same time bin between the telescope sites. Given that each shared time-bin contains a state of mean photon number $\epsilon$, the logical gate error rate needs to be much smaller than $\epsilon$. Per time bin, we need at least
    1. (a)
      One single-qubit gate to absorb the photon into the memory
    2. (b)
      One set of (logical) multi-qubit gates to encode the memory state onto a quantum error correcting code
    3. (c)
      One set of (logical) gates to perform parity checks
    4. (d)
      One set of logical single-qubit gates to perform the measurements. This one is non-Clifford in general
Below, we outline these core requirements and the current state of the art.
V.1 Phase stabilisation
Maintaining phase coherence across distant telescope sites is critical for interferometry. The entire system must be stabilised to within a small fraction of the optical wavelength. A shared optical phase reference is necessary. Notably, recent twin-field QKD experiments have achieved 98% interference visibility over 200 km at 1550 nm [36], suggesting that phase stabilisation over VLBI-scale baselines is technically feasible.
V.2 Time-bin multiplexing
Figure 4: The coherence time of the received signal depends inversely on the bandwidth. Signals with a path length difference larger than the coherence time will not interfere with one another. In this illustration, there is one photon shared non-locally between the two stations in bin $t_{1}$, and in $t_{5}$: the photon in $t_{1}$ will not interfere with the one in $t_{5}$.
Starlight must be parsed into discrete time bins, with each bin corresponding to the coherence time of the incoming thermal light. The coherence time of the light is inversely proportional to the bandwidth $\Delta\nu$ [29],
$\Delta t\sim\frac{1}{\Delta\nu}.$ (19)
As the bandwidth increases, the coherence time—and therefore the required time-bin size—decreases proportionally. As an example, for light generated by thermal sources with a narrow spectral width ($\Delta\nu\sim 10^{8}$ Hz), the corresponding coherence time is of order $10^{-8}s$, and coherence length is of order 3 m. For ease of comparison, Table. 5 provides a comparison for the frequency vs wavelength bandwidth and their conversion for various wavelengths.
Current state-of-the-art optical switching operates at rates of up to 38 MHz with approximately 80% transmission efficiency [37], corresponding to a minimum time-bin size of roughly $2.6\times 10^{-8}$ seconds. In contrast, at a wavelength of 1550 nm, a spectral bandwidth of 1 nm corresponds to a temporal coherence time of about 8 picoseconds—nearly 3000 times shorter than the achievable switching interval. This mismatch between the coherence times and switching capabilities is a challenge for time-bin multiplexing at optical frequencies.
One possible strategy to alleviate this constraint is to implement frequency multiplexing, dividing the incoming bandwidth into, for example, 100 discrete spectral channels. This effectively increases the usable time-bin size per channel, easing the demands on switching speed. Combined with recent advances in fast optical switches—particularly those based on barium titanate (BTO) platforms developed by groups such as PsiQuantum—this suggests that an order-of-magnitude improvement in switching speed is realistic in the near term. Such improvements would significantly narrow the gap between current switching capabilities and the time-resolution requirements of optical VLBI, making it more viable across a broad range of spectral bands.
V.3 Implementing quantum nodes with optical cavities
Figure 5: Schematic of a quantum node architecture with memory and communication qubits overcoupled to the communication bus — a fibre — enhanced by an optical cavity.
We would like to make use of the strong light-matter interaction enabled by state-of-the-art optical cavities to coherently transfer information from an external optical field into the stable states of an atom. In order to do this efficiently, it is necessary to implement a unitary transformation of all of the incident modes to the target cavity mode, before the light is coupled into memory qubits in a quantum node. Toward this task, it has been shown [38] that by exploiting the spatial dependence of impedance mismatch in a pair of misaligned Fabry-Perot resonators, high efficiency spatial mode conversion of optical photons is possible. It was demonstrated that light from a Hermite-Gauss HG00 mode can be converted into an arbitrary target HGm0 mode with conversion/transmission efficiency $>75\%$, by varying the length of a Fabry-Perot resonator over a few nanometers.
The many platforms for implementing quantum nodes in physical systems are built around memory qubits, which store the information in long-lived nuclear spin, dark or metastable states, perfectly isolated from the environment. To control and read out that memory, quantum nodes would typically use a direct, controlled coupling to a cavity, or mediate it via a short-lived communication qubit [39]. The cavity is used to boost the intrinsically low efficiency of these processes and enhance the in- and out-coupling from the memory to the fibre.
To this end, the cavity operates in the overcoupled regime, meaning that its coupling to the fiber mode $\kappa_{\text{out}}$ should dominate over other losses $\kappa_{0}$, and support high cooperativity $C=g^{2}/\kappa\gamma\gg 1$, defined by the cavity-qubit coupling $g$, and decoherence rate of the qubit $\gamma$.
High cooperativity serves to accelerate a particular, desired emission pathway for the qubit (i.e. into the fibre-coupled optical mode, thus reducing the role of other emission and dephasing channels), and boost the efficiency of in-coupling the incident photon into the qubit toward values comparable to $C/(1+C)$. Since the cooperativity can be conveniently approximated by the ratio of the quality factor and the effective volume of the mode $C\propto Q/(V/\lambda^{3})$ (with the latter normalised by the diffraction-limited volume $\lambda^{3}$), we can maximise it by embracing macroscopic high-$Q$ cavities (like the Fabry-Perot resonators), microscopic systems operating near the $V/\lambda^{3}\sim 1$ point (fiber cavities), or deeply sub-wavelength nanophotonic devices (photonic crystal cavities) [39, 40].
The particular choice of the platform for a quantum node is dictated by the realisation of the memory qubits, for example:
V.3.1 Atomic defects, incorporating nitrogen, silicon, and germanium atoms
Atomic defects, incorporating nitrogen, silicon, or germanium atoms next to vacancies in diamond (NV, SiV, and GeV) can implement both the communication and memory qubits, encoding information in the electronic transitions of the defect, nuclear nuclear spins of the nearby atoms—13C for NVs, and 29Si for SiVs—respectively [41, 42]. These two types of defects also offer very different challenges and opportunities: NV exhibits far longer coherence times at room temperature ($\sim 1$ s, several orders of magnitude over SiV), but also is far more sensitive to electrical noise and coupling to phonons, which degrades its optical response. Both defects can couple to moderate-$Q$ subwavelength nanoscale photonic crystal cavities, fabricated in diamond itself.
An example quantum network architecture was recently realised by Knaut et al. [43], with two nodes built around SiV in diamond photonic crystal cavities, operating below 200 mK. Memory qubits were realised using nuclear spin of a nearby 29Si (with 10 ms decoherence times, controlled by RF electronics), coupled to the fiber modes via a communications qubit (electronic spin states of SiV) and a cavity with moderate quality factor $Q\sim 10^{3}$ and cooperativity $C\sim 10$. Even accounting for the additional inefficiencies, including cavity-fiber coupling, correcting the frequency mismatching between the SiVs in the two nodes, downconverting optical emission from the SiV at 737 nm closer to the telecom wavelength at 1350 nm, and losses suffered by the photonic qubits over 45 km propagation between the nodes, the this simple network realised entanglement between the two memories with fidelity of 0.7 at 1 Hz rate.
With each mode built around micron-scale cavities, moderate requirements for cooling, and rapid improvements in techniques for diamond nanofabrications, these compact designs constitute a promising platform for implementing large-scale quantum nodes.
V.3.2 Neutral atoms and ions
Neutral atoms and ions, trapped (optically or electrically) in ultra-high vacuum, and coupled to high-$Q$ optical cavities [44]; thanks to efficient isolation from the environment, such memory qubits can exhibit coherence times approaching 1 s (for neutral atoms [45]) over 1 h (for ions [46]). Entanglement between remote nodes was demonstrated with 0.85 fidelity has been demonstrated with neutral atoms over a decade ago [47], and — more recently — over 0.9 with trapped ions [48].
A realisation of a quantum repeater, using two 87Rb atoms implementing memory qubits in a high-$Q$ cavity was reported by Langenfeld et al.  [49]. The decoherence rate of the qubits exceeded 20 ms, and the interfacing with optical photons was implemented via a Raman process, enhanced by a mm-length cavity with $Q\sim 10^{8}$, with the resulting cooperativity $C\approx 4$.
While trapped atoms are arguably the most should mature technology for realising quantum nodes, the enormous footprint of ultra-high vacuum and optical setups significantly limits their scalability, and opportunities for deployment in large-scale quantum networks.
V.4 Light-matter coupling gates
This light-matter described in the previous section must preserve the photon’s temporal and phase information, as these encode the interferometric signal of interest. Moreover, because each photon may only occupy a narrow time bin and arrive probabilistically, the entire absorption-and-storage operation must be repeatable and low-latency, with errors significantly below the mean photon number $\epsilon$ per bin. Any imperfection at this stage reduces the signal-to-noise ratio of the subsequent quantum measurement, and hence the overall sensitivity of the VLBI array.
The state-of-the-art gate fidelity and speeds for various platforms include:
  • ion traps: Single-qubit gates, error $1.5\times 10^{-6}$ per Clifford and gate time 600 ns [50], and Multi-qubit gates: Two-qubit gates at 270 $\mu s$ [51].
  • neutral-atoms : two-qubit gates via Rydberg blockade with error $5\times 10^{-3}$ and gate time $250$ns [52]; Ref. [53] reports a two-qubit logic gate using Förster resonances with error $0.01$ and gate time $6.5$ns. Single-site addressable measurements have been performed with an error of $7.3\times 10^{-3}$ over a time $36$ms [54].
  • Superconducting qubits (Google) [55]: $\sim$ 20 ns for single qubits and $\sim$ 40 ns for two-qubit gates.
The problem of storing and processing quantum information converted from optical photons into matter qubits for the purposes of quantum sensing does not require full fault tolerant quantum computing. This is because the logical operations are Clifford operations in which case it suffices to have elementary gate errors below the threshold for fault tolerant quantum error correction.
V.5 Entanglement distribution and repeaters
In quantum-enabled VLBI, pre-shared entanglement replaces the direct optical link between telescope sites, enabling nonlocal measurements without transmitting fragile stellar photons across long baselines. This requires distributing high-fidelity Bell pairs between nodes at rates compatible with the time-bin structure and detection probability. The quality and rate of entanglement distribution set key limits on the overall system performance.
Some state-of-the-art entanglement generation specifications include:
  • Experimental entanglement generation: by bulk-optics and polarisation-entangled photons, achieved key rate greater than 1 Gbit/sec at telecom [56]
  • In an integrated SiC platform, entangled photon pair generation rates $\sim 10^{4}$ pairs/s [57].
  • Commercial Entanglement Sources: Commercial products, such as those from Quantum Computing Inc., offer entangled photon sources with generation rates of 5M pairs/$\mu$W/s.
  • Photonic Crystal Cavities: Utilising photonic crystal cavities, time-energy entangled photon pairs have been generated at on-chip rates up to 22 MHz using only 36 $\mu$W of pump power.  [58]
In terms of repeaters, the latest progress includes:
[…]
  • Experimental demonstration of heralded entanglement between two absorptive quantum memories separated by 3.5 meters. This setup, utilizing polarisation-entangled photon-pair sources and solid-state quantum memories with up to 1 GHz bandwidth, successfully distributed maximally entangled states with a fidelity of 80.4 [61].
  • Towards a frequency-multiplexed repeater [62]: spectrally Multiplexed Photon-Pair Source; signal at 795 nm and idler at 1532 nm (telecom): up to 200 spectral modes, each mode $\sim$ 1.5 GHz (at 1532 nm), with 6.5 GHz spacing.
  • Proposal: Deterministic Quantum Repeater with Single Atoms in Cavities
Over the past decade, entanglement distribution rates have improved dramatically, driven by breakthroughs in source quality, detection efficiency, and the emergence of quantum memories and spectral/temporal multiplexing techniques. This marks an orders-of-magnitude increase in capability, especially when comparing today’s multiplexed and memory-integrated systems to earlier linear optics-only approaches.
These developments represent real progress. Demonstrations like the memory-enhanced Bell-state measurements using solid-state spins in nanophotonic cavities, or the use of spectrally multiplexed SPDC sources with quantum-memory-compatible bandwidths, have shown that entanglement generation rates can be significantly boosted, even under realistic loss conditions.
That said, we’re still far from practical, high-rate quantum repeater networks.
  • Rates remain low when memory is involved
  • Multiplexing is not fully integrated yet—many setups demonstrate multiplexed sources or filters, but don’t yet combine this with active feed-forward control and working quantum memories.
  • Most systems often require frequency conversion, which adds loss and complexity.
  • Scalability is an open challenge—interfacing multiple quantum memories, handling timing jitter, and maintaining fidelity across many links will require engineering solutions beyond current lab-scale demonstrations.
Below, we summarise the various protocols in the current literature.
VI Literature survey
Several protocols have made advances to quantum-enabled VLBI, we summarise a selection of the following protocols, which are listed in approximate chronological order of appearance.
  • Quantum Nonlocality in Weak-Thermal-Light Interferometry [1].
  • Longer-Baseline Telescopes Using Quantum Repeaters [2].
  • Quantum memories and the double-slit experiment: implications for astronomical interferometry [63].
  • Experiment demonstration [64] of the protocol in Ref. [2]
  • Optical Interferometry with Quantum Networks: scheme based on efficient quantum data compression offers an exponential decrease in required entanglement bandwidth [3].
  • Quantum-assisted telescope arrays: frequency and multi-domain extension to [3], GHZ states used for parity checks [4]
  • Using STIRP to absorb the stellar photons’ information into the quantum memories, and incorporate quantum error correction [5].
  • Optimal qubit circuits for quantum-enhanced telescopes [65]: encodes the time of arrival of the stellar photon into a quantum memory; reduces both the number of ancilla qubits and the number of gate operations by a factor of two.
  • Vacuum beam guide for large-scale quantum networks: remove most of the quantum tech and put a variable delay line back in [6].
  • Continuous variable teleportation schemes based on two-mode squeezed-vacuum [66, 67]
  • Comparison between amplitude and intensity interferometry as a function of baseline and mean photon number [68]
  • One experimental demonstration of intensity interferometry [69]
  • Non-local and phase reference resource states [70]
VI.1 Quantum Nonlocality in Weak-Thermal-Light Interferometry [1].
In Ref. [1], Tsang showed that, for any local measurement strategies, the Fisher information scales as $\epsilon^{2}$. Non-local measurements are required if we wish to achieve an FI linear in $\epsilon$. In particular, If we project the stellar state onto the basis |$\ket{\psi_{\delta}^{\pm}}\bra{\psi_{\delta}^{\pm}}$ |, then the FI matrix for direct detection is
$\bm{\mathrm{I}}=\frac{\epsilon}{1-\text{Re}(\gamma e^{-i\delta})^{2}}\left(\begin{array}[]{cc}\cos^{2}\delta&\sin\delta\cos\delta\\ \sin\delta\cos\delta&\sin^{2}\delta\\ \end{array}\right)$ (22)
with eigenvalues
$\lambda_{1}=0,\qquad\lambda_{2}=\frac{\epsilon}{1-\text{Re}(\gamma e^{-i\delta})^{2}}$ (23)
The trace norm of the FI matrix is then
$||\bm{\mathrm{I}}||=\lambda_{1}+\lambda_{2}\geq\epsilon$ (24)
Then for $M$ measurements,
$||\bm{\mathrm{I}}^{M}||=M||\bm{\mathrm{I}}||\geq M\epsilon$ (25)
Note that $||\bm{\mathrm{I}}^{M}||$ can be significantly larger than $\epsilon$ for $\gamma\rightarrow 1$, and that care should be taken when discussing the sum of the two terms: this is because per photon, $J_{\phi}$ is upper bounded by $1$, whereas $J_{\gamma}$ can, in principle, approach infinity.
VI.2 Longer-Baseline Telescopes Using Quantum Repeaters [2]
The problem with direct interferometry is that it is difficult to transport the single photon state over long distances without incurring loss. Ref. [2] proposes a protocol where, instead of sending a valuable quantum state directly over a noisy quantum channel, entanglement is distributed between the two telescopes instead.
Suppose we can distribute the following entangled state,
[…]
where $\delta$ is determined by a controllable phase, allowing completion of the protocol to determine $\phi$.
Now, each half shared state |$\ket{\psi_{\mathrm{shared}}}$ is then interfered with the stellar state at the respective telescope station with a 50:50 beam splitter (on each side, there are two detectors). The coincidence count can be used to directly directly to measure the $\gamma e^{i\phi}$. We post-select the measurement results, considering the events where a single photon is observed at $A$ and $B$ simultaneously. The total probability of seeing correlated photons at the output, conditioned on having one click at each telescope, is,
$\frac{1}{2}(1+\mathrm{Re}[\gamma e^{-i(\phi-\delta)}]),$ (27)
and the total probability of seeing an anti-correlation is,
$\frac{1}{2}(1-\mathrm{Re}[\gamma e^{-i(\phi+\delta)}]).$ (28)
This protocol requires that a “ground” photon interferes with a stellar photon, which means that the two need to be matched in their respective temporal and frequency modes. This implies that per photon measured, the entanglement consumption is of order $1/\epsilon$: see unary encoding below.
Ref. [71] extended the analysis to distributing more ground photons, where parallel-distributed ground photons interfere with the astronomical photon in a balanced multimode beam splitter. Ref. [72] investigated the detrimental effect of having distinguishable ground photons.
VI.3 Unary encoding
Let us now introduce quantum memory. Consider the simplest unary encoding where there is one memory qubit for each time bin. Given $M$ bins, after a single photon is stored, in the 5th time bin, for example, the memories will be in the state
$\frac{1}{\sqrt{2}}(\ket{000010...}_{A}\ket{000000...}+$
$e^{i\phi}\ket{000000...}_{A}\ket{000010...}_{B})$ (29)
Note that this is the same as
[…]
where the entangled state is the 5th qubits of systems $A$ and $B$.
Since the stellar state is mostly vacuum, we must figure out which memory qubits contain a stellar photon. This can be accomplished by using pre-shared Bell pairs and 2 CZ gates to perform a parity check, projecting out the vacuum [3] Denote the preshared Bell pair as |$\ket{\Phi}$,
[…]
$\epsilon\left(\frac{1+\gamma}{2}\right)\ket{\psi_{+}^{\phi}}\bra{\psi_{+}^{\phi}}\otimes\ket{\Phi^{-}}\bra{\Phi^{-}}+$
$\epsilon\left(\frac{1-\gamma}{2}\right)\ket{\psi_{-}^{\phi}}\bra{\psi_{-}^{\phi}}\otimes\ket{\Phi^{-}}\bra{\Phi^{-}}.$ | . (31)
That is, if a photon is present in the memory, a parity check leads to |$\ket{\Phi}$ acquiring a minus phase.
In a naive encoding, the unary encoding requires one entangled pair for each time bin, thus representing a large consumption of entanglement. Performing the entanglement-enabled parity checks will result in the following
|$\ket{\Phi^{+}}^{\otimes M-1}\ket{\Phi^{-}}$ (32)
consuming $M$ Bell pairs.
VI.4 Unary encoding with binary search
Figure 6: Schematic for binary search for the time bin into which the photon arrived. This protocol is a new result.
We can use an encoding that is efficient in the entanglement consumption, where the encoding is akin to a binary search. We choose a block size $M$ where $M\epsilon\sim 1$.
  • For each run of the parity check, we divide the memory block into two (Fig. 6).
  • To search for where the photon is within the time bin, we apply CZ gates between all of Alice’s or Bob’s memory qubits.
In our example with 4 time bins, a photon arrives in the second bin and is shared non-locally between Alice and Bob. The state is
$\rho\propto$ |$\ket{0,0}_{A_{1}B_{1}}(\ket{01}_{A_{2}B_{2}}\pm e^{i\phi}\ket{10}_{A_{2}B_{2}})$
|$\ket{0,0}_{A_{3}B_{3}}\ket{0,0}_{A_{4}B_{4}}$ (33)
We divide the block of 4 into two halves. The first half is
$\rho_{1,2}\propto$ |$\ket{0,0}_{A_{1}B_{1}}(\ket{01}_{A_{2}B_{2}}\pm e^{i\phi}\ket{10}_{A_{2}B_{2}})$ (34)
To perform the check, for all qubits in the half-block, CZ gates are applied between the memory qubit and the shared Bell pair.
$\rho_{1,2}\otimes\ket{\Phi^{+}}$ $\stackrel{{\scriptstyle 4\times\text{CZ}}}{{\rightarrow}}\rho_{1,2}\times\ket{\Phi^{-}}$ (35)
If there is no photon in the block, such as bins 3 and 4, then the Bell state remains the same.
$\rho_{3,4}\otimes\ket{\Phi^{+}}$ $\stackrel{{\scriptstyle 4\times\text{CZ}}}{{\rightarrow}}\rho_{3,4}\times\ket{\Phi^{+}}$ (36)
The procedure continues by iteratively dividing the block containing the photon into halves, until a single time bin remains.
This encoding is still linear in the memory requirements, where the number of memory qubits scales as $1/\epsilon$. However, the entanglement consumption has been reduced to scaling with $\sim~{}\log_{2}(1/\epsilon)$
VI.5 Binary encoding: optical interferometry with quantum Networks [3]
The binary encoding proposed in [3] is efficient in the entanglement consumption in the parity check. Let us
  • Assume approximately one photon arrives over $M$ time bins, which $M\sim 1/\epsilon$.
  • Label each time bin $m\in\mathbb{Z}_{+}$, with its binary representation $m_{2}$.
  • Define logical qubits |$\ket{\bar{0}}\equiv\ket{0...0}$ and |$\ket{\bar{1}_{m}}\equiv\ket{m_{2}}$. E.g. the fifth time bin
    |$\ket{\bar{1}_{5}}\equiv\ket{1010...0}$
A encoded logical CNOT is performed between the photonic degrees of freedom and the memory qubits at each time bin
|$\ket{0}(\ket{\bar{0}},\ket{\bar{1}_{j}})$ $\stackrel{{\scriptstyle\bar{\text{CX}_{m}}}}{{\rightarrow}}\ket{0}(\ket{\bar{0}},\ket{\bar{1}_{j}})$ (37)
|$\ket{1}(\ket{\bar{0}},\ket{\bar{1}_{j}})$ $\stackrel{{\scriptstyle\bar{\text{CX}_{m}}}}{{\rightarrow}}\ket{1}(\ket{\bar{1}_{m}},\ket{\bar{1}_{j}+\bar{1}_{m}})$ (38)
This encoding keeps track of the time-of-arrival information of the photon: empty time bins leave the memories unchanged.
$\bar{\rho}_{0}\equiv\ket{\bar{0},\bar{0}}\bra{\bar{0},\bar{0}}$ | (39)
Performing the CX operation,
$\rho_{AB}\otimes\bar{\rho}_{0}$ $\stackrel{{\scriptstyle\bar{\text{CX}_{m}}}}{{\rightarrow}}(1-\epsilon)\rho_{\text{vac}}\otimes\bar{\rho}_{0}$
$+\frac{\epsilon(1+|g|)}{2}\ket{\Psi^{+}_{\phi,m}}\bra{\Psi^{+}_{\phi,m}}$ |
$+\frac{\epsilon(1-|g|)}{2}\ket{\Psi^{-}_{\phi,m}}\bra{\Psi^{-}_{\phi,m}}$ | (40)
where
|$\ket{\Psi^{\pm}_{\phi,m}}=(\ket{0,1}_{A,B}\ket{\bar{0},\bar{1}_{m}}\pm e^{i\phi}\ket{1,0}_{A,B}\ket{\bar{1}_{m},\bar{0}})$ (41)
Then the photonic degree of freedom is decoupled from the memory qubit by measuring it in the X basis. (This step is likely challenging, which is solved in the next paper). Then, applying the appropriate correction (or post-processing), the memory qubits end up being
$\rho_{\star}=(1-\epsilon)\bar{\rho}_{0}+$ $\frac{\epsilon(1+|g|)}{2}\ket{\psi^{+}_{\phi,m}}\bra{\psi^{+}_{\phi,m}}$ |
$+$ $\frac{\epsilon(1-|g|)}{2}\ket{\psi^{-}_{\phi,m}}\bra{\psi^{-}_{\phi,m}}$ | (42)
where |$\ket{\psi^{\pm}_{\phi,m}}=\frac{1}{\sqrt{2}}(\ket{\bar{0},\bar{1}_{m}}\pm e^{i\phi}\ket{\bar{1}_{m},\bar{0}})$
In summary, if a photon arrives in the $(m-1)$-th time bin, then it is stored in such a way that the time bin is flipped according to |$\ket{\bar{1}_{m}}$.
This protocol uses the same set of memory for each time bin, which means that gates are applied to configure each subsequent time bin encoding. Given the short coherence time for each time bin, the gate speed needs to catch up with this operation.
VI.5.1 Error due to multiphoton contribution
Let us examine errors due to multiphoton contribution and how to alleviate them.
Starlight is thermal, and the block size is chosen such that $M\epsilon\sim 1$. If the average photon number in the block is 1, then the probability of having more than 1 photon in the block is also high. Let us examine the state post-encoding; using the notation of Ref. [3]
$\rho_{AB}=p_{0}\rho_{\text{vac}}+p_{1}\rho_{1}+(1-p_{0}-p_{1})\rho_{\text{multi}}$ (43)
where $p_{0}$ is the probability of receiving the vacuum, $p_{1}$ is the probability of having a single photon; and $\rho_{\text{multi}}$ is the state of the memory where more than one photon has landed. In the case of $\rho_{\text{multi}}$, the memory qubit is considered completely depolarised.
For $\gamma=1$, we have [3],
$p_{0}=\frac{1}{\epsilon+1},\qquad p_{1}=\frac{\epsilon}{(1+\epsilon)^{2}}$ (44)
If we have $\rho_{AB}^{\otimes M}$ in a block of $M$, then final decoded state in the readout qubit has coefficients given by a trinomial distribution of the input coefficients:
$\rho_{AB}^{\otimes M}\rightarrow$ $\left(\frac{1}{1+\epsilon}\right)^{M}\rho_{\text{vac}}+M\left(\frac{1}{1+\epsilon}\right)^{M-1}\left(\frac{\epsilon}{(1+\epsilon)^{2}}\right)\rho_{1}$
$+\left(1-\frac{1}{(1+\epsilon)^{M}}-\frac{M\epsilon}{(1+\epsilon)^{M+1}}\right)\rho_{\text{multi}}$ (45)
Figure 7: Showing $c$ in Eq. (46) as a function of $\epsilon$. The value on the y-axis is $c$, for (top) $M\epsilon=1$, (bottom) $M\epsilon=0.1$. The expense is that the entanglement consumption rate went up by a factor of 10.
Assuming the multiphoton state is completely mixed, the post-selected ‘single photon’ state reads as
$\rho=c\rho_{1}+(1-c)\rho_{\text{mix}},\qquad c=\frac{M\epsilon}{[(1+\epsilon)^{M}-1](1+\epsilon)}$ (46)
This presents a problem, because for high-contrast $c$ needs to be as close to 1 as possible, and this is fundamental; $c$ depends on how we choose $M\epsilon$ (Fig. 7).
For this protocol, the number of required memory qubits is approximately $\log_{2}(1/\epsilon)$ [4]. Let us calculate the distributed entanglement (|$\ket{\Phi}$) consumption rate:
  1. 1.
    The time bin size (coherence time) is given by the inverse of the frequency bandwidth, $\tau=1/\Delta\nu$. Therefore, there will be $\Delta\nu$ time bins per second.
  2. 2.
    The average number of photons per second is $\Delta\nu\epsilon$.
  3. 3.
    Per stellar photon, the parity check requires $\log_{2}(1/\epsilon)$; we arrive at the entanglement consumption rate being approximately $\Delta\nu\times\epsilon\log_{2}(1/\epsilon)$.
  4. 4.
    We likely need to increase this by a factor of approximately 10 to circumvent multiphoton events (see Fig. 7). The multi-photon problem may be lifted if we expand to broadband operation (section below).
For $\Delta\nu$ = 10 GHz, $\epsilon=7\times 10^{-7}$, entanglement distribution rate is $\Delta\nu\epsilon\log_{2}(1/\epsilon)\sim$ 200 kHz. See Table 3 for further examples.
$\lambda$ $\Delta\nu$ (Hz) $\Delta\lambda$ $\epsilon$ $\Delta\nu\epsilon\log_{2}(1/\epsilon)$ per sec
555 nm 10 GHz $10^{-11}$m $10^{-7}$ $23\times 10^{3}$
555 nm 10 GHz $10^{-11}$m $10^{-10}$ 33
555 nm 1 THz 1 nm $10^{-7}$ $2\times 10^{6}$
555 nm 1 THz 1 nm $10^{-10}$ $3\times 10^{3}$
555 nm 1 THz 1 nm $10^{-11}$ 40
760 nm 500 GHz 1 nm $10^{-7}$ $1.2\times 10^{6}$
760 nm 500 GHz 1 nm $10^{-10}$ $1.2\times 10^{3}$
760 nm 500 GHz 1 nm $10^{-12}$ 20
1.65 $\mu$m 110 GHz 1 nm $10^{-7}$ $2.5\times 10^{5}$
1.65 $\mu$m 110 GHz 1 nm $10^{-10}$ $3.6\times 10^{2}$
1.65 $\mu$m 110 GHz 1 nm $10^{-12}$ 4.4
Table 3: Entanglement consumption rate for various wavelengths, bandwidths, and mean photon numbers.
VI.6 Quantum-assisted telescope arrays [4]
The above protocol is generalised to broadband operation in Ref [4] The incoming light is split into R frequency bands. As an example, consider that a photon arrives in the 5th time bin, and is a mixture of two frequency bands.
$\rho\approx(1-\epsilon)\rho_{\text{vac}}+\frac{\epsilon}{2}(\rho_{\nu 1}+\rho_{\nu 2})$ (47)
$\rho_{\nu 1}=$ $|\displaystyle\frac{1}{2}(\ket{0,1}\bra{0,1}+\ket{1,0}\bra{1,0}+g_{1}\ket{0,1}\bra{1,0}$ |
$+$ $g_{1}^{*}\ket{1,0}\bra{0,1})_{A,B,\nu 1}\otimes\ket{0,0}\bra{0,0}_{A,B,\nu 2}$
$\rho_{\nu 2}=$ |$\ket{0,0}\bra{0,0}_{A,B,\nu 1}\otimes\frac{1}{2}(\ket{0,1}\bra{0,1}+$
[…]
The notation here is:
|$\ket{\underbrace{101}_{\text{time}}\underbrace{1}_{\text{freq}},0000}$ (49)
The first 3 memory qubits encode the time bin, and the 4th encodes the frequency band.
In this particular example, integrating the stellar photon into the memory leads to the following transformation.
|$\ket{0,0}_{\nu 1}\ket{0,0}_{\nu 2}\ket{0000_{A},0000_{B}}$
$\rightarrow$ |$\ket{0,0}_{\nu 1}\ket{0,0}_{\nu 2}\ket{0000_{A},0000_{B}},$ (50)
|$\ket{1,0}_{\nu 1}\ket{0,0}_{\nu 2}\ket{0000_{A},0000_{B}}$
$\rightarrow$ |$\ket{1,0}_{\nu 1}\ket{0,0}_{\nu 2}\ket{1010_{A},0000_{B}},$ (51)
|$\ket{0,1}_{\nu 1}\ket{0,0}_{\nu 2}\ket{0000_{A},0000_{B}}$
$\rightarrow$ |$\ket{0,0}_{\nu 1}\ket{0,1}_{\nu 2}\ket{0000_{A},1010_{B}},$ (52)
|$\ket{0,0}_{\nu 1}\ket{1,0}_{\nu 2}\ket{0000_{A},0000_{B}}$
$\rightarrow$ |$\ket{0,0}_{\nu 1}\ket{1,0}_{\nu 2}\ket{1011_{A},0000_{B}},$ (53)
|$\ket{0,0}_{\nu 1}\ket{0,1}_{\nu 2}\ket{0000_{A},0000_{B}}$
$\rightarrow$ |$\ket{0,0}_{\nu 1}\ket{0,1}_{\nu 2}\ket{0000_{A},1011_{B}}.$ (54)
This variation has parallel operation over frequencies, at the expense of memory scaling as $R\log_{2}M$, but the entanglement consumption scales logarithmically with $R$ and $M$.
One technical difficulty associated with using the time bin encoding is that we need to encode the memory qubits as fast as the time bins. This timescale is extremely fast compared to the current gate times for most platforms.
VI.7 Incorporating quantum error correction [5]
Figure 8: Cavity-assisted coherent single-photon transfer. (a) A system of qubits has logical states |$\ket{0_{L}},\ket{1_{L}}$; ancillary qubit 1 is initially prepared into a Bell state with the register, $1/\sqrt{2}(\ket{0_{L}0_{G}}+\ket{1_{L}1_{G}})$ ; ancilla 2 is used in the STIRAP interaction to interact with the star photon. (b) Energy levels of ancilla atom 2 used for the STIRAP interaction.
We can incorporate quantum error correction (QEC) as follows. We assume the stellar photon is described by the state in Eq. (IV).
In this protocol, the set-up is depicted in Fig. 8: (a) inside a cavity, we use three different sets of systems. We denote the blue array as the register. The blue array is initialised in a codespace of a QEC code encoding a single logical qubit, spanned by its logical codewords |$\ket{0_{L}}$ and |$\ket{1_{L}}$. We also need ancilla qubit 1 (green), and ancilla atom 2 (red). Note that the three types of matter qubits could consist of different electronic sublevels of the same species of atom if desired.
Suppose we now prepare the register and the green ancilla (here the subscript $G$ denotes green) in the Bell state
[…]
$\frac{1}{\sqrt{2}}\left(\ket{1_{R},0_{R}}_{AB}\pm e^{i\phi}\ket{0_{R},1_{R}}_{AB}\right).$ (56)
Performing a Bell measurement on the red and green ancillae teleports the state onto the registers. After the Pauli operator correction dependent on the measurement outcome, the state of the registers between Alice and Bob becomes an entangled state, and the entanglement arises entirely from the starlight photon.
Since the initial starlight state is mixed, after the encoding, the density matrix shared between Alice and Bob is
$\rho_{AB}\approx$ $(1-\epsilon)\ket{0_{L}0_{L}}\bra{0_{L}0_{L}}_{AB}+\epsilon\left(\frac{1+\gamma}{2}\right)\ket{\psi_{+,L}^{\phi}}\bra{\psi_{+,L}^{\phi}}$ |
$+\epsilon\left(\frac{1-\gamma}{2}\right)\ket{\psi_{-,L}^{\phi}}\bra{\psi_{-,L}^{\phi}}+O(\epsilon^{2}),$ (57)
where |$\ket{\psi^{\phi}_{\pm,L}}=(\ket{0_{L},1_{L}}\pm e^{i\phi}\ket{1_{L},0_{L}})/\sqrt{2}.$
Figure 9: Population transfer for of a three-state adiabatic passage. Top: the interaction strengths of $g$ and $\Omega$ as a function of time $t$ in units of $1/g$. Bottom: occupancy of in |$\ket{e},\ket{0}_{R}$ and |$\ket{1}_{R}$ in ancillary atom 2 (r). The detuning parameter is set to $\Delta(t)=g^{2}+\Omega^{2}(t)$ to satisfy the adiabatic condition [73]. Note that the excited state |$\ket{e}$ is not populated.
The technically difficult part of this protocol includes:
  • The STIRAP step needs to be extremely low noise: the stellar state is unencoded at this point, and the error rate must be $\ll\epsilon$.
  • For a complete population transfer, here $t=50$, which appears long. The cavity quality needs to be sufficiently high to retain the stellar photon until the transfer is complete.
  • STIRAP schemes have very narrow bandwidth. (Rare earth: 10k Hz)
As an example of a suitable STIRAP scheme, consider a 87Rb atom trapped in an optical tweezer and coupled to a fibre Fabry-Perot cavity as recently demonstrated [74]. Adapted to the goal here of coherent single-photon transfer, we follow the prescription in [75]. Choose qubit states |$\ket{0}=\ket{5^{2}S_{1/2},F=1,m_{F}=0}$, |$\ket{1}=\ket{5^{2}S_{1/2},F=2,m_{F}=0}$, and |$\ket{e}=\ket{5^{2}P_{3/2},F=3,m_{F}=0}$, where the linewidth of the |$\ket{1}\leftrightarrow\ket{e}$ transition ($\lambda=780\,$nm) is $\gamma=2\pi\times 6\,$MHz (FWHM). We assume a cavity finesse $F\approx 2\times 10^{5}$, a waist radius ${\rm w}_{r}\approx 2\,\mu$m and a length $L\approx 40\,\mu$m resulting in a cooperativity of $C=3\lambda^{2}F/(2\pi^{3}{\rm w}_{r}^{2})\approx 1500$ (as demonstrated in [76]) with a coupling strength of $g=\sqrt{3\lambda^{2}C\gamma}/(2\pi^{2}{\rm w}_{r}^{2}L)\approx 2\pi\times 400\,$MHz and $\kappa=\pi C/LF\approx 2\pi\times 20\,$MHz (FWHM), so that $\gamma/\kappa\approx 0.3$. Over a population transfer period of $T$ one estimates a fidelity of $F=e^{-T\kappa}$ which for $T=50g^{-1}$ would be too small, $F<0.1$. Better cavities providing a factor of $10$ improvement in $g/\kappa$ would provide $F>0.75$. Note Tweezer induced dephasing rates on state |$\ket{1}$ can be treated as negligible $\gamma_{\phi}/g=0.03\times 10^{-6}$ [77].
VI.8 Continuous-variable investigations
States received from astronomical sources are inherently thermal [29]. Therefore, it is natural to consider this problem in the framework of continuous-variable (CV) quantum information [78, 79].
In Ref. [66] three schemes are scrutinised: (1) direct interferometry (DI), where the two modes of the stellar state are physically brought together for interference. Each mode suffers transmission loss parameterised by $\eta\in[0,1]$. (2) A local strategy, where heterodyne detection is performed separately on the modes held by Alice and Bob, and no loss is incurred. (3) A CV teleportation-based [80] strategy, where a two-mode squeezed vacuum (TMSV) is distributed to Alice and Bob. During distribution, each mode of the TMSV suffers transmission loss parameterised by $\eta$. Bob performs joint homodyne measurements as prescribed by standard CV teleportation and sends his measurement outcomes to Alice.
In the lossless regime, it is shown that a squeezing parameter of $r\approx 2$ (18 dB) is required to reach $\approx$ 95% of the QFI achievable with DI; such a squeezing level is beyond what has been achieved experimentally. In the low-loss regime, the CV teleportation strategy becomes inferior to DI, and the performance gap widens as loss increases. Curiously, in the high-loss regime, a small region of loss exists where the CV teleportation strategy slightly outperforms both DI and local heterodyne, representing a transition in the optimal strategy.
However, the advantage is very limited because it occurs for a small region of loss, and the magnitude of the advantage is also small. Practical difficulties further impede achieving any quantum advantage, limiting the merits of a CV teleportation-based strategy for stellar interferometry. Conclusion: TMSV teleportation-based strategies are practically infeasible.
VI.9 Vacccum beam guide for quantum networks [6]
This paper introduces a simple and elegant solution for long-distance distribution of quantum states: a ground-based optical channel composed of a vacuum tube with precisely aligned lenses. The authors demonstrate that a VBG with realistic parameters can achieve an attenuation rate as low as $3\times 10^{-5}$ dB/km. The system avoids the complexity of quantum repeaters by requiring only passive optical elements. Key sources of loss—lens imperfections, residual gas absorption, and beam misalignment—are rigorously analysed and shown to be manageable with current technology. The system avoids the complexity of quantum repeaters by requiring only passive optical elements. Key sources of loss—lens imperfections, residual gas absorption, and beam misalignment—are rigorously analysed and shown to be manageable with current technology.
This architecture is not only promising for quantum communication but also enables transformative advances in quantum-enabled long-baseline optical interferometry (QVLBI). In particular, we see two distinct, yet complementary, directions in which the VBG infrastructure could be adapted for interferometric science.
VI.9.1 High-Efficiency Entanglement Distribution for Quantum Interferometry
In a repeater-based VLBI architecture, VBGs could serve as the physical backbone for entanglement distribution between telescope nodes. This offers several key advantages:
  • Low-loss transmission over thousands of kilometres enables high-rate distribution of entangled photons with reduced need for quantum repeaters.
  • The VBG channel can support a variety of quantum encodings (e.g., time-bin, polarisation, frequency-bin) suitable for distributed quantum metrology and interferometry.
VI.9.2 Using the VBG as an interferometer arm
Alternatively, the VBG itself can be directly repurposed as an interferometric guide channel, effectively acting as one arm of a giant optical interferometer:
In this approach, starlight collected at separate telescopes is coupled into the VBG and transported to a common detection point for interference-based measurement. Since the VBG avoids the material losses and dispersion of optical fibres, it provides an ideal medium for coherent light propagation over long baselines.
However, in this configuration, one must reintroduce active variable path compensation to maintain fringe stability within a fraction of an optical wavelength. This could be achieved through delay lines and phase-locking mechanisms, possibly guided by a bright reference source or classical beacon.
This direct-interference approach would closely mimic classical optical interferometry but at scales (1000+ km) previously accessible only in the radio domain.
In addition, if implemented on the Moon, the vacuum beam guide would benefit from the natural hard vacuum of the lunar environment, eliminating the need for complex evacuated tubes. This could significantly reduce engineering overhead while enabling ultra-long baseline quantum interferometry with minimal optical loss.
VII Open questions
Implementing quantum-enabled VLBI is extremely technologically demanding: we need to resolve (the many) individual time modes corresponding to sparsely arriving stellar photons. This requires highly efficient coupling into quantum memories with mode-selective control. The large amount of time and frequency modes make this a difficult task.
There are open questions that remain. First, what is the best way to transfer the information in a single photon into a quantum memory? What are the trade-offs between adiabatic loading schemes (e.g. STIRAP) and faster dynamical techniques? Adiabatic loading schemes (e.g. STIRAP) offer robustness but are slow, while dynamical schemes allow faster coupling at the cost of greater sensitivity to timing and control. The optimal choice depends on memory coherence time and system timing constraints.
How many temporal modes do we actually need? Does this depend solely on $\epsilon$, or do detector and memory limitations impose stricter bounds? What should the mode shapes look like in practice? For binary encoding, the number of memory qubits scales as $\log(1/\epsilon)$, but practical constraints such as detector dead time and memory reset times also play a role.
Is temporal or wavelength multiplexing unavoidable? If so, how can we implement it without overwhelming the memory and feedforward control? Can we design architectures that minimise multiplexing while preserving sensitivity? While multiplexing boosts photon capture and entanglement rates, it increases system complexity. Minimising multiplexing while maintaining detection efficiency remains an open design question.
In summary, mode-resolved quantum VLBI imposes a unique set of challenges that bridge quantum optics, temporal mode engineering, and real-time feedback. Solving these problems will be essential for unlocking the full potential of quantum-enhanced astronomical imaging and precision geodesy.
VIII Concluding remarks
This report has outlined the feasibility and potential of quantum-enabled optical Very Long Baseline Interferometry (VLBI). By integrating quantum information protocols, many long-standing challenges in optical VLBI can be mitigated.
From a technological perspective, we examined a range of architectures, including repeater-based entanglement distribution, binary-encoded quantum memories, and hybrid protocols. We believe the uses of the vacuum beam guide architecture, either as a lossless medium for entanglement distribution or as a coherent optical delay channel for direct interferometry, offer scalable paths toward kilometre-scale optical baselines.
Encouragingly, most of the individual hardware components—such as high-cooperativity optical cavities, long-lived quantum memories, and entangled photon sources—are now approaching the threshold of practical feasibility. Experimental progress in photon collection, phase stabilisation, and entanglement distribution suggests that several essential subsystems could be prototyped with near-term technologies.
The primary bottlenecks that remain are the speed and fidelity of quantum gate operations. In particular, time-bin multiplexed architectures impose stringent requirements on gate execution times and error rates that exceed current capabilities in most physical platforms. These limitations currently constrain the implementation of real-time encoding, decoding, and error correction procedures necessary for robust operation.
Nevertheless, the convergence of quantum technologies toward these benchmarks is rapid. Continued improvements in quantum control and photonic integration may soon enable the deployment of quantum-enhanced VLBI systems capable of achieving diffraction-limited resolution at optical wavelengths. Such a capability would unlock new frontiers in both astronomical imaging and precision Earth science.
Acknowledgements.
We thank Max Charles, John Bartholomew, Michael Steel, Christian Schwab, John Jeffers, Lorenzo Maccone, and Chiara Macchiavello for insightful discussions. ZH is supported by an ARC DECRA Fellowship (DE230100144) “Quantum-enabled super-resolution imaging” and the EPSRC International Network in Space Quantum Technologies (EP/W027011/1). MKS is supported by an Australian Research Council Discovery Early Career Researcher Award (DE220101272). DO is supported by the EPSRC Quantum Technology Hub in Quantum Communication (EP/T001011/1), EPSRC Integrated Quantum Networks Research Hub (EP/Z533208/1), and ESA (4000147561/25/NL/FGL/ss). PK is supported by EPSRC’s Large Baseline Quantum-Enhanced Imaging Networks grant (EP/V021303/1) and Mathematical Tools for Practical Quantum Imaging Protocols grant (UKRI069). We acknowledge the traditional custodians of the Macquarie University land, the Wallumattagal clan of the Dharug nation, whose cultures and customs have nurtured and continue to nurture this land since the Dreamtime. We pay our respects to Elders past, present and future.
Appendix A Tables for reference
Wavelength $\lambda$ (nm) Frequency $\nu$ (THz) $\Delta\lambda$ (pm)
600 500 0.12
1500 200 0.75
10000 30 33
Table 4: Conversion of a frequency bandwidth of $\Delta\nu=10^{8}$ Hz to wavelength bandwidth at different central wavelengths. The coherence time is $10$ ns, as per Eq. (19).
Wavelength $\lambda$ (nm) Frequency $\nu$ (THz) $\Delta\nu$ (THz)
400 750 18.8
500 600 12.0
600 500 8.3
1000 300 3.0
1550 193.5 1.25
2000 150 0.75
Table 5: Conversion of a 10 nm wavelength bandwidth to frequency bandwidth at different central wavelengths.
Appendix B Calculating the mean photon numbers of nearby stars and exoplanets
Planet mass ($M_{\oplus}$) temperature (K) distance (light year) host apparant brightness ($m_{AB}$)
Earth 1 255 0 N/A
Proxima Centauri b $\geq 1.07$ 228 4.2 11.05
Ross 128 b $\geq 1.4$ 280 11 11.13
GJ 1061 d $\geq 1.64$ 218 12 13.03
Luyten b $\geq 2.89$ 258 12.4 9.85
Gliese 1002 b $\geq 1.08$ 231 15.8 13.837
Table 6: A list of nearby exoplanets with potentially habitable temperature, their masses relative to Earth, distance, and magnitudes of their host star
In this section, we include details of how we arrive at the mean photon numbers of nearby stars and exoplanets. Here the monochromatic absolute magnitude $m_{AB}$ is defined as the logarithm of a spectral flux density with the usual scaling of astronomical magnitudes The zero-point is approximately about 3631 Jy, where
[…]
Denote $F_{\nu}$ as the spectral flux density
$m_{\text{AB}}$ $\approx-2.5\log_{10}\left(\frac{F_{\nu}}{3631~{}\text{Jy}}\right)$ (59)
The relationship between $m_{AB}$ and flux is
$F_{\nu}=10^{-0.4(m_{AB}+48.6)}\text{erg}~{}s^{-1}\text{cm}^{-2}\text{Hz}^{-1}$ (60)
Convert this to SI units
[…]
$1~{}\text{cm}^{2}$ $=10^{-4}\text{m}^{2}$ (62)
Therefore in SI units, $F_{\nu}$ has units of $10^{-3}$ W m-2 Hz-1:
$F_{\nu}=10^{-0.4(m_{AB}+48.6)}\times 10^{-3}\text{W m}^{-2}\text{Hz}^{-1}$ (63)
Converting to flux per unit wavelength $F_{\lambda}$
$F_{\lambda}=\frac{F_{\nu}\times c}{\lambda^{2}}$ (64)
where $c$ is the speed of light and $\lambda$ is the wavelength. Calculate the photon flux density
$\Phi_{\lambda}=\frac{F_{\lambda}}{hc/\lambda}=\frac{F_{\nu}}{h\times\lambda}$ (65)
Now, to calculate the total photon specific bandwidth, we assume the density is approximately constant over the bandwidth (this holds for narrow bandwidths):
$\Phi_{\text{tot}}$ $=\Phi_{\lambda}\times\Delta\lambda$ (66)
$=\frac{1}{h\times\lambda}10^{-0.4(m_{AB}+48.6)}\times 10^{-3}\times\Delta\lambda$ (67)
Now, to relate that to $\epsilon$, we need to figure out ’how long’ the photon is, i.e. the coherence time. Then we’ll just call it the size of the time bin, here $\Delta\nu$ is the frequency bandwidth
$\tau_{c}=\frac{1}{\Delta\nu}$ (68)
For a 10 GHZ bandwidth signal, $\Delta\tau\sim 10^{-10}$.
To convert between a frequency bandwidth to a wavelength bandwidth,
$\Delta\lambda=\Delta\nu\times\lambda^{2}/c$ (69)
$\Delta\nu=\Delta\lambda\times c/\lambda^{2}$ (70)
Repeating the calculation of [PRL 123, 070504 (2019)]: for a $\Delta\nu=10$ GHz, a total collection area of $10^{2}$,
[…]
For the CHARA array, visible interferometry has bandwith between 10-50 nm.
Putting in the numbers, $\Delta\nu=10\times 10^{9}$, $\lambda=555\times 10^{-9},m=10$
$\epsilon$ $=\Phi_{\text{tot}}\times\tau_{c}\times 10$ (72)
$\approx 10^{-7}$ (73)
In comparison, [4] obtains $\epsilon\approx 7\times 10^{-7}$.
Now, let us examine at the bands of interest, corresponding to bio-signature molecular absorption bands:
Host ($m_{\text{AB}}$) $\lambda$ $\Delta\lambda$ $\Delta\nu$ Photons/s ($\Phi_{\text{tot}}\times$ area) $\tau_{c}$ $\epsilon$ (star) $\epsilon$ (planet)
9 760 nm 1 nm 500 GHz 181106 $\times 10^{-12}$ $3.5\times 10^{-7}$ $3.5\times 10^{-16}$
9 1.65 $\mu$m 1 nm 100 GHz 83418 $10^{-12}$ $7.5\times 10^{-7}$ $7.5\times 10^{-16}$
11 760 nm 10 nm 5 THz 287035 2×$\times 10^{-13}$ $5.5\times 10^{-8}$ $5.5\times 10^{-17}$
11 1.65 $\mu$m 10 nm 1.1 THz 132210 $10^{-12}$ $1.2\times 10^{-7}$ $1.2\times 10^{-16}$
11 760 nm 1 nm 500 GHz 28703 $\times 10^{-12}$ $5.5\times 10^{-8}$ $5.5\times 10^{-17}$
11 1.65 $\mu$m 1 nm 100 GHz 13221 $10^{-11}$ $1.2\times 10^{-7}$ $1.2\times 10^{-16}$
13 760 nm 1 nm 500 GHz 4549 $\times 10^{-12}$ $9\times 10^{-9}$ $9\times 10^{-18}$
13 1.65 $\mu$m 1 nm 100 GHz 2095 $10^{-12}$ $1.9\times 10^{-8}$ $1.9\times 10^{-17}$
Table 7: Expected number of photons to collect at given wavelength, bandwidth for the host star. The wavelengths of interest are: 760nm is the absorption line for molecular oxygen; for Earth-like conditions the pressure-broadening linewidth is typically $\sim$ 2 pm at FWHM; 1.65 $\mu$m is a methane absorption line, and for earth-like conditions the FWHM ×$\times 10^{-4}\mu$m. Here we have assumed a collection area of 10 m2. Spectral demultiplexing doesn’t decrease $\epsilon$ due to the fact that that the effective time bins get larger. We assume a conservative $10^{-9}$ factor between the star and planet’s brightness. These numbers might differ depending on what assumption is made about the temperature of the star.

### v2

@@
 Abstract
-Optical Very Long Baseline Interferometry offers the potential for unprecedented angular resolution in both astronomical imaging and precision measurements.
-Classical approaches, however, face significant limitations due to photon loss, background noise, and the requirements for dynamical delay lines over large distances.
-This document surveys recent developments in quantum-enabled VLBI, which aim to address these challenges using entanglement-assisted protocols, quantum memory storage, and nonlocal measurement techniques.
+Optical Very Long Baseline Interferometry (VLBI) offers the potential for unprecedented angular resolution in both astronomical imaging and geodesy measurements.
+Classical approaches face limitations due to photon loss, background noise, and their need for dynamical delay lines over large distances.
+This document surveys recent developments in quantum-enabled optical VLBI, which aim to address these challenges using entanglement-assisted protocols, quantum memory storage, and nonlocal measurement techniques.
 While its application to astronomy is well known, we also examine how these techniques may be extended to geodesy – specifically, the monitoring of Earth’s rotation.
 Particular attention is given to quantum-enhanced telescope architectures, including repeater-based long-baseline interferometry and quantum error-corrected encoding schemes, which offer a pathway toward high-fidelity optical VLBI.
-To aid the discussion, we also compare specifications for key enabling technologies to current state-of-the-art experimental components, including switching rates, gate times, entanglement distribution rates, and memory lifetimes.
+To aid the discussion, we also compare specifications for key enabling technologies to current state-of-the-art experimental components.
 By integrating quantum technologies, future interferometric networks may achieve diffraction-limited imaging at optical and near-infrared wavelengths, surpassing the constraints of classical techniques and enabling new precision tests of astrophysical and fundamental physics phenomena.
@@
 This is especially true for light originating in astronomical objects, where we cannot illuminate the objects of interest, and we are restricted to analysing the light that reaches us.
-Several challenges hinder the progress in building large-baseline optical interferometers, including the weight of the instrument itself, photon loss, background noise, and the requirements for long dynamical delay lines.
-These factors ultimately limit the achievable distance between telescope sites.
-Quantum technologies can help bypass transmission losses: using quantum memories and entanglement, we can replace the direct optical links, allowing for much larger distances.
-Among the key scientific applications of optical Very Large Baseline Interferometry (VLBI), two stand out.
-First, VLBI offers a path toward high angular resolution astronomical imaging, particularly in regimes that are inaccessible to conventional telescopes.
-Second, and less commonly considered, optical VLBI can achieve unprecedented spatial and temporal resolution in geodetic metrology, such as monitoring Earth’s rotation, tectonic motion, and lunar libration.
-Today’s most advanced interferometric imaging systems operate at radio and microwave wavelengths.
-Instruments such as the Event Horizon Telescope (EHT) achieve resolutions of tens of microarcseconds by combining a maximum baseline of $\sim$12,000 km (roughly the Earth’s diameter) with a typical wavelength of 3 mm.
-Shifting to optical wavelengths can, in principle, increase resolution by several orders of magnitude.
-For instance, an optical system operating at $\sim$600 nm would outperform the EHT in resolution with a baseline as short as 2.4 km.
-At 1550 nm (a standard C-band telecom wavelength), the equivalent threshold is 6.2 km.
-II Astronomical imaging
 2019 saw a dramatic demonstration of the image resolution as limited by the size of the aperture of the imaging system.
@@
 The data is then post-processed to reconstruct an image.
-Moving to optical frequencies, however, introduces significant complications: even the fastest electronics cannot directly measure the oscillations of the electric field at optical frequencies.
-Nevertheless, if these complications can be overcome, the potential gains in resolution are substantial.
-Optical VLBI offers a 3-5 orders of magnitude improvement in resolution over radio-based systems, depending on the wavelength and baseline.
-This opens the door to imaging astrophysical structures that are beyond the reach of our current imaging systems.
-Optical VLBI refers to interferometric regimes where real-time calibration of optical path lengths is not feasible, necessitating the use of closure phases and post-processed visibilities [7].
+However, direct measurement is not possible in optics, because even the fastest electronics cannot directly measure the oscillations of the electric field at optical frequencies.
+Nevertheless, if we can find alternate solutions, we can potentially gain 3-5 orders of magnitude improvement in resolution.
+This unlocks possibilities for imaging astrophysical structures that are beyond the reach of our current imaging systems.
+Several challenges hinder the progress in building large-baseline optical interferometers, including the weight of the instrument itself, photon loss, background noise, and the requirements for long dynamical delay lines.
+These factors ultimately limit the achievable distance between telescope sites.
+Quantum technologies can help bypass transmission losses: using quantum memories and entanglement, we can replace the direct optical links, allowing for much larger distances.
+If deployed on the Moon, we can exploit the natural vacuum as a beam guide for long-baseline operation without atmospheric distortion.
+Among the key scientific applications of optical VLBI, two stand out.
+First, VLBI offers a path toward high angular resolution astronomical imaging.
+Second, and less commonly considered, optical VLBI can achieve unprecedented spatial and temporal resolution in geodetic metrology, such as monitoring Earth’s rotation, tectonic motion, and lunar libration.
+Optical VLBI refers to interferometric regimes where real-time compensation of optical path lengths is not feasible, necessitating the use of closure phases and post-processed visibilities [7].
 Instruments like CHARA [8] and VLTI [9] operate in this regime, and future concepts such as the Big Fringe Telescope [10] aim to extend it.
-Our work does not propose optical VLBI as a new idea, but rather explores how quantum-enhanced protocols might overcome key technical limitations, particularly in photon loss and scalability.
+Our work does not propose optical VLBI as a new idea, but rather explores how quantum-enabled protocols might overcome key technical limitations, particularly in photon loss and scalability.
+These approaches are complementary to other classical strategies such as heterodyne interferometry [11], and intensity interferometry [12].
+However, the quantum approach uniquely targets large-baseline, high-precision measurements and scalable architecture without the need for formation flying or complex delay-line infrastructure.
+Today’s most advanced interferometric imaging systems operate at radio and microwave wavelengths.
+Instruments such as the Event Horizon Telescope (EHT) achieve resolutions of tens of microarcseconds by combining a maximum baseline of $\sim$12,000 km (roughly the Earth’s diameter) with a typical wavelength of 3 mm.
+Shifting to optical wavelengths can potentially increase resolution by several orders of magnitude.
+An optical system operating at 600 nm could match the EHT with only a 2.4 km baseline; at 1550 nm, 6.2 km would suffice.
+II Astronomical imaging
 The advantage of a quantum approach is not merely improved angular resolution based on extending the baseline, but achieving the best precision in complex visibility measurements.
 Many science cases—such as detecting exoplanets, imaging stellar surfaces, or tracking relativistic orbital precession—rely on high signal-to-noise measurements rather than resolving power alone.
-Current optical interferometers face severe constraints on scalability and baseline length, due to the need for physical delay lines and lossy multi-way beam combination.
-Quantum architectures offer a route to nonlocal beam combination by connecting telescopes without direct optical interference.
-If deployed on the Moon, we can exploit the natural vacuum as a beam guide for long-baseline operation without atmospheric distortion.
-These approaches are complementary to other classical strategies such as heterodyne interferometry [11], and intensity interferometry [12].
-However, the quantum approach uniquely targets large-baseline, high-precision measurements and scalable architecture without the need for formation flying or complex delay-line infrastructure.
-In the remainder of this section discuss some science cases for using large-baseline optical interferometry, including imaging of exoplanets, tracking the stars around black holes, and measuring separations of double-star systems.
+This section discusses some science cases for using large-baseline optical interferometry, including imaging of exoplanets, tracking the stars around black holes, and measuring separations of double-star systems.
 II.0.1 Imaging exoplanets
@@
 Table 1: Angular separation of nearby exoplanets from their parent stars.
-Exoplanets present a formidable observational challenge: they are both extremely small in angular size and orders of magnitude fainter than their host stars.
-Direct imaging of these objects requires two key capabilities:
-  • High angular resolution, required to spatially separate the planet from the glare of the star.
-  • High visibility (contrast), required to detect the planet’s faint signal against the much brighter stellar companion.
+Imaging exoplanets is a challenging task: they are both extremely small in angular size and orders of magnitude fainter than their host stars. Direct imaging of these objects requires:
+  • high angular resolution to spatially separate the planet from the glare of the star.
+  • high visibility (contrast) to detect the planet’s faint signal against the much brighter parent star.
 Optical VLBI offers a potential solution to both challenges.
@@
 At this resolution, it becomes conceivable to directly image spatially resolved features in the atmospheres of nearby exoplanets, opening new possibilities for studying their composition, structure, and potentially even weather systems.
-We show the separation of the star from the exoplanet in Table 1.
+We show the separation of the star from the exoplanet in Table 1 for reference.
 II.0.2 Tracking stars orbiting black holes
@@
 The GRAVITY collaboration (see, e.g., Refs.
-[14, 15, 16]) currently achieves near-infrared resolution of 3 milli-arcseconds and astrometric precision of 10–100 $\upmu$as, and a baseline of 130 m.
+[14, 15, 16]) currently achieves a near-infrared resolution of 3 milli-arcseconds and astrometric precision of 10–100 $\upmu$as, and a baseline of 130 m.
 To detect the Lense–Thirring precession at comparable wavelengths (near infrared), a baseline of at least $130\times 240=32$ km would be required.
-Figure 1: The Lense–Thirring effect: a star’s orbit precesses due to the frame-dragging caused by a rotating massive object, such as a spinning black hole.
+Figure 1: The Thirring effect: a star’s orbit precesses due to the frame-dragging caused by a rotating massive object, such as a spinning black hole.
 II.0.3 Measuring the stellar separation of A. Centauri.
@@
 III Geodesy
-Figure 2: Geometry for using VLBI for geodesy. The position of the reference star is assumed fixed, and the goal is to measure the baseline vector $\bm{B}$.
+Figure 2: Geometry for using VLBI for geodesy. The position of the reference star is assumed to be fixed, and the goal is to measure the baseline vector $\bm{B}$.
+Geodesy uses precise measurements of Earth’s shape, orientation, and gravity field—often through VLBI and satellite techniques—to monitor tectonic motion, the Earth’s rotation, and maintain global reference frames.
 VLBI is a geometric technique; it measures the time difference between the arrival at two Earth-based antennas emitted by a distant quasar.
@@
 Relative changes in the antenna locations from a series of measurements indicate tectonic plate motion, regional deformation, and local uplift or subsidence.
-The core idea is the precise timing of light signals:
-  1. 1.
 When a quasar emits EM waves, they arrive at different telescopes at slightly different times due to Earth’s shape and rotation.
-  2. 2.
 The exact time delay between telescopes depends on the Earth’s rotation state at that moment.
-  3. 3.
-    By comparing the signals and modelling their delay based on Earth’s motion, scientists can infer tiny variations in rotation speed and axis position.
-Geodetic VLBI plays a crucial role in geodesy, enabling precise measurements of the Earth’s radio telescope positions and long-term motion of the tectonic plates.
-Currently, the tectonic motion is measured with accuracy up to 0.1 mm/year, though the seasonal variations of the individual radio telescopes’ positions are also monitored [17].
+By comparing the signals and modelling their delay based on Earth’s motion, we can infer small variations in rotation speed and axis position.
+Geodetic VLBI plays a crucial role in geodesy.
+Currently, the tectonic motion is measured with an accuracy up to 0.1 mm/year, though the seasonal variations of the individual radio telescopes’ positions are also monitored [17].
 Some of the radio telescopes were affected by strong earthquakes (Japan, 2010; Chile, 2011) [18].
@@
 In the 1990s, the leap second was introduced almost every year.
-Although the Earth rotation deceleration nowadays is not so high and the leap second is introduced rarely (the last time was in 2015), the situation may change again and needs to be monitored continually [19].
+Although the Earth’s rotation deceleration nowadays is not so high and the leap second is introduced rarely (the last time was in 2015), the situation may change again and needs to be monitored continually [19].
 Leap seconds introduce significant challenges for computing, satellite navigation, and global timekeeping systems.
@@
 To mitigate these issues, discussions are ongoing regarding the potential elimination or redefinition of leap seconds in the future [20].
-Optical VLBI offers substantial improvement in the temporal resolution of UT1 measurements, enabling detection of changes in Earth’s rotation on shorter and more precise timescales.
+Optical VLBI offers a substantial improvement in the temporal resolution of UT1 measurements, enabling the detection of changes in Earth’s rotation on shorter and more precise timescales.
 III.1.2 Polar motion
@@
 VLBI helps in tracking polar motion, as the Earth’s rotational axis drifts slightly due to mass redistributions, such as the melting of ice sheets and shifts in the planet’s mantle [21][22].
-Optical VLBI can enhance our ability to track polar motion with finer spatial precision due to the shorter wavelengths involved, potentially improving sensitivity to small-scale mass redistributions.
+Optical VLBI can enhance our ability to track polar motion with finer spatial resolution, potentially improving sensitivity to small-scale mass redistributions.
 III.1.3 Precession and nutation
@@
 These catalogues are important for tracking deep space missions [28].
-If optical VLBI is used, the enhancement is a finer determination of source positions in the celestial reference frame, with potential uncertainties reduced by 3-5 orders of magnitude compared to radio VLBI.
 IV The model and parameter estimation
 High-resolution astronomical imaging and high-precision geodesy are fundamentally linked by their reliance on the same core technology: interferometry.
-In astronomical imaging, interferometers are used to resolve fine spatial structure by coherently combining light collected from a distant source.
-In geodesy, the principle is inverted: rather than using interferometry to study a celestial target, the celestial source serves as a fixed reference, and the interferometric measurement is used to track the motion of the Earth itself.
-In both contexts, the required technology is essentially the same, therefore, in this work, we adopt a common physical model for the stellar source.
-To facilitate analytic comparison and benchmarking, we focus on the well-understood two-mode model, for which the quantum Fisher information is readily computable and serves as a natural metric of the resolution.
+In astronomical imaging, interferometers are used to resolve fine spatial structures by coherently combining light collected from a distant source.
+In geodesy, the principle is inverted: the celestial source serves as a fixed reference, and the interferometric measurement is used to track the motion of the Earth itself.
+Given the shared technology, we adopt a common physical model for the stellar source and focus on the well-understood two-mode case, where quantum Fisher information serves as a natural metric for comparison.
 Consider a single frequency band whose two-mode continuous variable quantum description is given by a mean vector and covariance matrix of the form [29, 30, 5],
@@
 where we have used the quadrature ordering $(q_{A},p_{A},q_{B},p_{B})$, with subscripts referring to Alice ($A$) and Bob ($B$).
-In the limit that $\epsilon\ll 1$, which is typically true for astronomy, we can describe the optical state by the density matrix,
+In the limit that $\epsilon\ll 1$, which is typically true for astronomy, we can describe the state by the density matrix,
 $\rho_{\star}\approx$ $(1-\epsilon)\ket{\text{vac},\text{vac}}\bra{\text{vac},\text{vac}}_{AB}+$
@@
 […]
-If we were to divide a bandwidth spectrum (centred at 500 THz or 600 nm) of $\Delta\lambda=10$ nm into 100 MHz narrow bands, this is approximately 8300 bands; the photon will have a coherence time of approximately 10 ns with a coherence length 3 m; if instead we divide this into 1 GHz bands (830 bands), then the coherence time will be 1 ns, and the coherence length will be 30 cm.
-In the literature, the bandwidth of the stellar source is often quoted as either frequency or wavelength; we include Table 4 for reference and ease of converting between them.
 For imaging, the parameters of interest are $\phi$ and $\gamma$, where $\phi\in[0,2\pi)$ is related to the location of the sources, and $\gamma\in[0,1]$ is proportional to the Fourier transform of the intensity distribution via the van Cittert-Zernike theorem [29].
@@
 $(\Delta\theta)^{2}\geqslant\frac{1}{NJ_{\theta}(\hat{\rho}_{\theta})}\,,$ (12)
-where $N$ is the number of copies of $\hat{\rho}_{\theta}$ used and $J_{\theta}$ is the quantum Fisher information (QFI) associated with the state $\hat{\rho}_{\theta}$.
-If there are multiple parameters we want to estimate, where $\bm{\mathrm{\theta}}=(\theta_{1},\theta_{2},\dots)$, we can define a QFI matrix $\bm{\mathrm{J}}$ that quantifies not only the QFI for each parameter (diagonal components) but also for correlated parameters (off-diagonal components).
-The matrix elements are given by
+$N$ is the number of copies of $\hat{\rho}_{\theta}$ used and $J_{\theta}$ is the quantum Fisher information (QFI) associated with the state $\hat{\rho}_{\theta}$.
+If there are multiple parameters we want to estimate, where $\bm{\mathrm{\theta}}=(\theta_{1},\theta_{2},\dots)$, we can define a QFI matrix $\bm{\mathrm{J}}$. The matrix elements are given by
 $J_{jk}:=\frac{1}{2}\text{Tr}[\hat{\rho}_{\bm{\mathrm{\theta}}}(\hat{L}_{j}\hat{L}_{k}+\hat{L}_{k}\hat{L}_{j})],$ (13)
@@
 $\operatorname{Cov}(\bm{\mathrm{\theta}})\geq\frac{1}{N}\bm{\mathrm{J}}^{-1}.$ (14)
-The QFI matrix elements for the incoming stellar state are found to be
+The QFI matrix elements for the incoming stellar state are [36]
 $J_{\phi}$ $=\frac{2\gamma^{2}\epsilon}{2+\epsilon(1-\gamma^{2})},$ (15a)
 $J_{\gamma}$ $=\frac{2\epsilon\left(2+\epsilon+\epsilon\gamma^{2}\right)}{\left(1-\gamma^{2}\right)\left(4+4\epsilon+\epsilon^{2}\left(1-\gamma^{2}\right)\right)},$ (15b)
-$J_{\phi\gamma}$ $=0$ (15c)
-If we take the equivalent trace norm of the QFI matrix,
+$J_{\phi\gamma}$ $=0.$ (15c)
+If we take the trace norm of the QFI matrix,
 $||\mathbf{J}||=J_{\phi}+J_{\gamma}$
@@
 $=\frac{(1+\gamma^{2}-\gamma^{4})}{1-\gamma^{2}}\epsilon-\frac{1}{2}\left(1+\gamma^{2}-\gamma^{4}\right)\epsilon^{2}+O(\epsilon^{3})$ (16)
-Note that the above expression can be significantly larger than $\epsilon$ for $\gamma\rightarrow 1$.
-It’s also perhaps unfair to lump them together because $J_{\phi}/\epsilon$ is capped by $1$, whereas $J_{\gamma}/\epsilon$ can, in principle, approach infinity.
+Note that Eq.
+(IV) can be significantly larger than $\epsilon$, and that care should be taken to discuss the sum $J_{\phi}$ of $J_{\gamma}$, because because $J_{\phi}/\epsilon\leq 1$, whereas $J_{\gamma}/\epsilon$ can, in principle, approach infinity.
 For geodesy, the task is to measure the distance between two telescope sites, given that a stable astronomical source is at a known location (Fig.
@@
 The most accurate way of measuring this distance is by estimating the optical path length difference between the two sites, which manifests as a phase shift $\phi$.
-      Requirements       Direct detection       quantum-enabled
+      Requirements       Direct interferometry       quantum-enabled
       Phase reference and phase stabilisation
@@
 |$\ket{\psi}$ $=\frac{1}{\sqrt{2}}(\ket{01}+e^{i\phi}\ket{10}),$
-$\phi$ $=\omega\tau,$
-$=\frac{B\sin\theta}{\lambda}.$ (17)
-From quantum metrology, we know that the QFI of $\phi$ is equal to 1; i.e. $\Delta\phi\approx 1$. Here we assume $\theta$ to be fixed, and $\lambda$ is also fixed. This means
+$\phi$ $=\omega\tau=\frac{B\sin\theta}{\lambda}.$ (17)
+From quantum metrology theory, we know that the QFI of $\phi$ is equal to 1; i.e. $\Delta\phi\approx 1$. Assume $\theta$ and $\lambda$ to be fixed, this means
 $\Delta B\geq\frac{\sin\theta}{\lambda}\frac{\Delta\phi}{\sqrt{n}},$ (18)
 where $n$ is the number of photons received.
-The relation in Eq.
-(18) reveals a key advantage of operating at optical wavelengths: since the uncertainty in the baseline measurement scales inversely with $\lambda$, moving from radio to optical can potentially yield a substantial gain in precision.
-For the same number of detected photons and observational geometry, shorter wavelengths directly translate into increased precision of geodetic measurements.
-V Technological requirements of OVLBI
+Eq.
+(18) highlights a key advantage of optical wavelengths: because baseline uncertainty scales inversely with $\lambda$, shorter wavelengths significantly improve geodetic precision for the same photon count and geometry.
+V Technological requirements of Optical VLBI
 Realising quantum-enabled optical VLBI involves a multi-layered technological framework that integrates several key components (see Table 2):
@@
   2. 2.
-    Collect the starlight and time-multiplex them into different bins (see Fig. 4)
+    Starlight is collected and time-multiplexed into different bins (see Fig. 4)
   3. 3.
@@
 Non-local measurement is performed between quantum memories of the same time bin between the telescope sites.
-Given that each shared time-bin contains a state of mean photon number $\epsilon$, the logical gate error rate needs to be much smaller than $\epsilon$.
 Per time bin, we need at least
@@
     4. (d)
-      One set of logical single-qubit gates to perform the measurements. This one is non-Clifford in general
+      One set of logical single-qubit gates to perform the measurements. This one is non-Clifford in general.
+Given that each shared time-bin contains a state of mean photon number $\epsilon$, the logical gate error rate needs to be much smaller than $\epsilon$.
 Below, we outline these core requirements and the current state of the art.
@@
 Notably, recent twin-field QKD experiments have achieved 98% interference visibility over 200 km at 1550 nm [37], suggesting that phase stabilisation over VLBI-scale baselines is technically feasible.
+A quantitative analysis on the shared phase reference is given in Ref.
+[38].
 V.2 Time-bin multiplexing
@@
 In this illustration, there is one photon shared non-locally between the two stations in bin $t_{1}$, and in $t_{5}$: the photon in $t_{1}$ will not interfere with the one in $t_{5}$.
-Starlight must be parsed into discrete time bins, with each bin corresponding to the coherence time of the incoming thermal light.
+Starlight must be parsed into discrete time bins, with each bin corresponding to the coherence time of the incoming thermal light (see Fig 4), because signal separated by longer than the coherence time will not interfere .
 The coherence time of the light is inversely proportional to the bandwidth $\Delta\nu$ [29],
 $\Delta t\sim\frac{1}{\Delta\nu}.$ (19)
+The bandwidth of the stellar source is often quoted in either frequency or wavelength.
 As the bandwidth increases, the coherence time—and therefore the required time-bin size—decreases proportionally.
 As an example, for light generated by thermal sources with a narrow spectral width ($\Delta\nu\sim 10^{8}$ Hz), the corresponding coherence time is of order $10^{-8}s$, and coherence length is of order 3 m.
-For ease of comparison, Table.
-5 provides a comparison for the frequency vs wavelength bandwidth and their conversion for various wavelengths.
 Current state-of-the-art optical switching operates at rates of up to 38 MHz with approximately 80% transmission efficiency [39], corresponding to a minimum time-bin size of roughly $2.6\times 10^{-8}$ seconds.
 In contrast, at a wavelength of 1550 nm, a spectral bandwidth of 1 nm corresponds to a temporal coherence time of about 8 picoseconds—nearly 3000 times shorter than the achievable switching interval.
-This mismatch between the coherence times and switching capabilities is a challenge for time-bin multiplexing at optical frequencies.
 One possible strategy to alleviate this constraint is to implement frequency multiplexing, dividing the incoming bandwidth into, for example, 100 discrete spectral channels.
@@
 To control and read out that memory, quantum nodes would typically use a direct, controlled coupling to a cavity, or mediate it via a short-lived communication qubit [41].
-The cavity is used to boost the intrinsically low efficiency of these processes and enhance the in- and out-coupling from the memory to the fibre.
-To this end, the cavity operates in the overcoupled regime, meaning that its coupling to the fiber mode $\kappa_{\text{out}}$ should dominate over other losses $\kappa_{0}$, and support high cooperativity $C=g^{2}/\kappa\gamma\gg 1$, defined by the cavity-qubit coupling $g$, and decoherence rate of the qubit $\gamma$.
+The cavity is used to boost the intrinsically low efficiency of these processes and enhance the in and out-coupling from the memory to the fibre.
+To this end, the cavity operates in the over coupled regime, meaning that its coupling to the fibre mode $\kappa_{\text{out}}$ should dominate over other losses $\kappa_{0}$, and support high cooperativity $C=g^{2}/\kappa\gamma\gg 1$, defined by the cavity-qubit coupling $g$, and decoherence rate of the qubit $\gamma$.
 High cooperativity serves to accelerate a particular, desired emission pathway for the qubit (i.e.
@@
 V.3.1 Atomic defects, incorporating nitrogen, silicon, and germanium atoms
-Atomic defects, incorporating nitrogen, silicon, or germanium atoms next to vacancies in diamond (NV, SiV, and GeV) can implement both the communication and memory qubits, encoding information in the electronic transitions of the defect, nuclear nuclear spins of the nearby atoms—13C for NVs, and 29Si for SiVs—respectively [41, 42].
+Atomic defects, incorporating nitrogen, silicon, or germanium atoms next to vacancies in diamond (NV, SiV, and GeV) can implement both the communication and memory qubits, encoding information in the electronic transitions of the defect, nuclear spins of the nearby atoms—13C for NVs, and 29Si for SiVs—respectively [43, 44].
 These two types of defects also offer very different challenges and opportunities: NV exhibits far longer coherence times at room temperature ($\sim 1$ s, several orders of magnitude over SiV), but also is far more sensitive to electrical noise and coupling to phonons, which degrades its optical response.
@@
 [45], with two nodes built around SiV in diamond photonic crystal cavities, operating below 200 mK.
-Memory qubits were realised using nuclear spin of a nearby 29Si (with 10 ms decoherence times, controlled by RF electronics), coupled to the fiber modes via a communications qubit (electronic spin states of SiV) and a cavity with moderate quality factor $Q\sim 10^{3}$ and cooperativity $C\sim 10$.
-Even accounting for the additional inefficiencies, including cavity-fiber coupling, correcting the frequency mismatching between the SiVs in the two nodes, downconverting optical emission from the SiV at 737 nm closer to the telecom wavelength at 1350 nm, and losses suffered by the photonic qubits over 45 km propagation between the nodes, the this simple network realised entanglement between the two memories with fidelity of 0.7 at 1 Hz rate.
+Memory qubits were realised using the nuclear spin of a nearby 29Si (with 10 ms decoherence times, controlled by RF electronics), coupled to the fiber modes via a communications qubit (electronic spin states of SiV) and a cavity with moderate quality factor $Q\sim 10^{3}$ and cooperativity $C\sim 10$.
+Even accounting for the additional inefficiencies, including cavity-fibre coupling, correcting the frequency mismatching between the SiVs in the two nodes, downconverting optical emission from the SiV at 737 nm closer to the telecom wavelength at 1350 nm, and losses suffered by the photonic qubits over 45 km propagation between the nodes, the this simple network realised entanglement between the two memories with fidelity of 0.7 at 1 Hz rate.
 With each mode built around micron-scale cavities, moderate requirements for cooling, and rapid improvements in techniques for diamond nanofabrications, these compact designs constitute a promising platform for implementing large-scale quantum nodes.
@@
 The decoherence rate of the qubits exceeded 20 ms, and the interfacing with optical photons was implemented via a Raman process, enhanced by a mm-length cavity with $Q\sim 10^{8}$, with the resulting cooperativity $C\approx 4$.
-While trapped atoms are arguably the most should mature technology for realising quantum nodes, the enormous footprint of ultra-high vacuum and optical setups significantly limits their scalability, and opportunities for deployment in large-scale quantum networks.
-V.4 Light-matter coupling gates
-This light-matter described in the previous section must preserve the photon’s temporal and phase information, as these encode the interferometric signal of interest.
+While trapped atoms are arguably the most mature technology for realising quantum nodes, the enormous footprint of ultra-high vacuum and optical setups significantly limits their scalability, and opportunities for deployment in large-scale quantum networks.
+V.4 Quantum gates
+This light matter described in the previous section must preserve the photon’s temporal and phase information, as these encode the interferometric signal of interest.
 Moreover, because each photon may only occupy a narrow time bin and arrive probabilistically, the entire absorption-and-storage operation must be repeatable and low-latency, with errors significantly below the mean photon number $\epsilon$ per bin.
 Any imperfection at this stage reduces the signal-to-noise ratio of the subsequent quantum measurement, and hence the overall sensitivity of the VLBI array.
-The state-of-the-art gate fidelity and speeds for various platforms include:
-  • ion traps: Single-qubit gates, error $1.5\times 10^{-6}$ per Clifford and gate time 600 ns [50], and Multi-qubit gates: Two-qubit gates at 270 $\mu s$ [51].
-  • neutral-atoms : two-qubit gates via Rydberg blockade with error $5\times 10^{-3}$ and gate time $250$ns [52]; Ref.
-[53] reports a two-qubit logic gate using Förster resonances with error $0.01$ and gate time $6.5$ns.
-Single-site addressable measurements have been performed with an error of $7.3\times 10^{-3}$ over a time $36$ms [54].
-  • Superconducting qubits (Google) [55]: $\sim$ 20 ns for single qubits and $\sim$ 40 ns for two-qubit gates.
-The problem of storing and processing quantum information converted from optical photons into matter qubits for the purposes of quantum sensing does not require full fault tolerant quantum computing.
-This is because the logical operations are Clifford operations in which case it suffices to have elementary gate errors below the threshold for fault tolerant quantum error correction.
+Recent experimental advances have demonstrated high gate fidelity and speeds for various platforms include
+  • Ion traps: Single-qubit gates with error $1.5\times 10^{-6}$ and 600 ns gate time [52]; two-qubit gates in 270 $\mu$[53].
+  • Neutral atoms: Rydberg-blockade gates with $\sim 5\times 10^{-3}$ error in 250 ns [54]; Förster resonance gates achieving 1% error in 6.5 ns [55]; single-site readout errors $\sim 7\times 10^{-3}$ over 36 ms [56].
+  • Superconducting qubits: Single- and two-qubit gate times of 20 ns and 40 ns, respectively, as demonstrated by Google [57].
+Importantly, quantum sensing in this context does not require universal, fault-tolerant quantum computation.
+Maintaining gate errors below standard quantum error correction thresholds is sufficient for reliable operation.
 V.5 Entanglement distribution and repeaters
-In quantum-enabled VLBI, pre-shared entanglement replaces the direct optical link between telescope sites, enabling nonlocal measurements without transmitting fragile stellar photons across long baselines.
-This requires distributing high-fidelity Bell pairs between nodes at rates compatible with the time-bin structure and detection probability.
-The quality and rate of entanglement distribution set key limits on the overall system performance.
-Some state-of-the-art entanglement generation specifications include:
-  • Experimental entanglement generation: by bulk-optics and polarisation-entangled photons, achieved key rate greater than 1 Gbit/sec at telecom [56]
-  • In an integrated SiC platform, entangled photon pair generation rates $\sim 10^{4}$ pairs/s [57].
-  • Commercial Entanglement Sources: Commercial products, such as those from Quantum Computing Inc., offer entangled photon sources with generation rates of 5M pairs/$\mu$W/s.
-  • Photonic Crystal Cavities: Utilising photonic crystal cavities, time-energy entangled photon pairs have been generated at on-chip rates up to 22 MHz using only 36 $\mu$W of pump power.  [58]
-In terms of repeaters, the latest progress includes:
+In quantum-enabled VLBI, pre-shared entanglement substitutes for direct optical links between telescope sites, enabling nonlocal measurements without transmitting fragile stellar photons across long baselines.
+The performance of such a system is fundamentally limited by the rate and fidelity of Bell pair distribution.
+Recent progress in entangled photon sources includes polarization-entangled photons at telecom wavelengths have enabled key rates exceeding 1 Gbit/s [58]; in SiC systems, pair generation rates of $\sim 10^{4}$ pairs/s have been demonstrated[59]; commercial sources: off-the-shelf systems, such as those from Quantum Computing Inc., now reach 5M pairs/$\mu$W/s; on-chip generation of time-energy entangled pairs has achieved 22 MHz using only 36 $\mu$W of pump power  [60].
 […]
-  • Experimental demonstration of heralded entanglement between two absorptive quantum memories separated by 3.5 meters.
-This setup, utilizing polarisation-entangled photon-pair sources and solid-state quantum memories with up to 1 GHz bandwidth, successfully distributed maximally entangled states with a fidelity of 80.4 [61].
-  • Towards a frequency-multiplexed repeater [62]: spectrally Multiplexed Photon-Pair Source; signal at 795 nm and idler at 1532 nm (telecom): up to 200 spectral modes, each mode $\sim$ 1.5 GHz (at 1532 nm), with 6.5 GHz spacing.
-  • Proposal: Deterministic Quantum Repeater with Single Atoms in Cavities
-Over the past decade, entanglement distribution rates have improved dramatically, driven by breakthroughs in source quality, detection efficiency, and the emergence of quantum memories and spectral/temporal multiplexing techniques.
-This marks an orders-of-magnitude increase in capability, especially when comparing today’s multiplexed and memory-integrated systems to earlier linear optics-only approaches.
-These developments represent real progress.
-Demonstrations like the memory-enhanced Bell-state measurements using solid-state spins in nanophotonic cavities, or the use of spectrally multiplexed SPDC sources with quantum-memory-compatible bandwidths, have shown that entanglement generation rates can be significantly boosted, even under realistic loss conditions.
-That said, we’re still far from practical, high-rate quantum repeater networks.
-  • Rates remain low when memory is involved
-  • Multiplexing is not fully integrated yet—many setups demonstrate multiplexed sources or filters, but don’t yet combine this with active feed-forward control and working quantum memories.
-  • Most systems often require frequency conversion, which adds loss and complexity.
-  • Scalability is an open challenge—interfacing multiple quantum memories, handling timing jitter, and maintaining fidelity across many links will require engineering solutions beyond current lab-scale demonstrations.
-Below, we summarise the various protocols in the current literature.
+These developments mark a substantial progress from earlier linear-optics-only approaches, thanks to improvements in source quality, detection efficiency, and the emergence of quantum memories and spectral/temporal multiplexing techniques.
+However, significant challenges remain before practical, high-rate repeater networks are feasible:: rates remain low when memory is involved; multiplexing is not fully integrated yet-—many setups demonstrate multiplexed sources or filters, but do not combine this with active feed-forward control and working quantum memories; most systems often require frequency conversion, which adds loss and complexity; scalability is an open challenge—interfacing multiple quantum memories, handling timing jitter, and maintaining fidelity across many links will require engineering solutions beyond current lab-scale demonstrations.
 VI Literature survey
-Several protocols have made advances to quantum-enabled VLBI, we summarise a selection of the following protocols, which are listed in approximate chronological order of appearance.
+Many protocols have made advances to quantum-enabled VLBI, we summarise a selection of the following; these are listed in approximate chronological order.
   • Quantum Nonlocality in Weak-Thermal-Light Interferometry [1].
-  • Longer-Baseline Telescopes Using Quantum Repeaters [2].
-  • Quantum memories and the double-slit experiment: implications for astronomical interferometry [63].
-  • Experiment demonstration [64] of the protocol in Ref. [2]
-  • Optical Interferometry with Quantum Networks: scheme based on efficient quantum data compression offers an exponential decrease in required entanglement bandwidth [3].
-  • Quantum-assisted telescope arrays: frequency and multi-domain extension to [3], GHZ states used for parity checks [4]
-  • Using STIRP to absorb the stellar photons’ information into the quantum memories, and incorporate quantum error correction [5].
-  • Optimal qubit circuits for quantum-enhanced telescopes [65]: encodes the time of arrival of the stellar photon into a quantum memory; reduces both the number of ancilla qubits and the number of gate operations by a factor of two.
-  • Vacuum beam guide for large-scale quantum networks: remove most of the quantum tech and put a variable delay line back in [6].
-  • Continuous variable teleportation schemes based on two-mode squeezed-vacuum [66, 67]
-  • Comparison between amplitude and intensity interferometry as a function of baseline and mean photon number [68]
-  • One experimental demonstration of intensity interferometry [69]
+  • Longer-Baseline Telescopes Using Quantum Repeaters [2], and an experimental demonstration [66].
+  • Optical Interferometry with Quantum Networks [3].
+  • Quantum-assisted telescope arrays [4]: frequency and time-multiplexed extension to [3].
+  • Quantum memories and the double-slit experiment (hard drive-based): implications for astronomical interferometry [67].
+  • Incorporating quantum error correction [5].
+  • Optimal qubit circuits for quantum-enhanced telescopes [68].
+  • Vacuum beam guide for large-scale quantum networks [6].
+  • Continuous variable teleportation schemes based on two-mode squeezed-vacuum [36, 69]; and fundamental limits in the absence of non-Gaussian resources [70].
+  • Comparison between amplitude and intensity interferometry as a function of baseline and mean photon number [71], see also the comparison with “weak heterodyne scheme” [72].
+  • Experimental demonstration of intensity interferometry [73]
   • Non-local and phase reference resource states [38]
@@
 $||\bm{\mathrm{I}}^{M}||=M||\bm{\mathrm{I}}||\geq M\epsilon$ (25)
-Note that $||\bm{\mathrm{I}}^{M}||$ can be significantly larger than $\epsilon$ for $\gamma\rightarrow 1$, and that care should be taken when discussing the sum of the two terms: this is because per photon, $J_{\phi}$ is upper bounded by $1$, whereas $J_{\gamma}$ can, in principle, approach infinity.
+Note that $||\bm{\mathrm{I}}^{M}||$ can be significantly larger than $\epsilon$ for $\gamma\rightarrow 1$.
 VI.2 Longer-Baseline Telescopes Using Quantum Repeaters [2]
@@
 [2] proposes a protocol where, instead of sending a valuable quantum state directly over a noisy quantum channel, entanglement is distributed between the two telescopes instead.
-Suppose we can distribute the following entangled state,
+Suppose we can distribute the following entangled state (“ground” photon),
 […]
 where $\delta$ is determined by a controllable phase, allowing completion of the protocol to determine $\phi$.
-Now, each half shared state |$\ket{\psi_{\mathrm{shared}}}$ is then interfered with the stellar state at the respective telescope station with a 50:50 beam splitter (on each side, there are two detectors).
-The coincidence count can be used to directly directly to measure the $\gamma e^{i\phi}$.
-We post-select the measurement results, considering the events where a single photon is observed at $A$ and $B$ simultaneously.
-The total probability of seeing correlated photons at the output, conditioned on having one click at each telescope, is,
-$\frac{1}{2}(1+\mathrm{Re}[\gamma e^{-i(\phi-\delta)}]),$ (27)
-and the total probability of seeing an anti-correlation is,
+Now, each half of |$\ket{\psi_{\mathrm{shared}}}$ is then interfered with the stellar state in Eq.
+(IV) at the respective telescope station with a 50:50 beam splitter.
+Then, we post-select and consider the events where a two-photon coincidence is observed at stations $A$ and $B$ simultaneously.
+The total probabilities of seeing photons in correlated, and anticorrelated outputs are respectively
+$\frac{1}{2}(1+\mathrm{Re}[\gamma e^{-i(\phi-\delta)}]),$
 $\frac{1}{2}(1-\mathrm{Re}[\gamma e^{-i(\phi+\delta)}]).$ (27)
-This protocol requires that a “ground” photon interferes with a stellar photon, which means that the two need to be matched in their respective temporal and frequency modes.
-This implies that per photon measured, the entanglement consumption is of order $1/\epsilon$: see unary encoding below.
+This protocol requires the “ground” photon to to interfere with a stellar photon, which means that the two need to be matched in their respective temporal and frequency modes.
+This implies that per stellar photon measured, the entanglement consumption is of order $1/\epsilon$: see unary encoding below.
 Ref.
-[71] extended the analysis to distributing more ground photons, where parallel-distributed ground photons interfere with the astronomical photon in a balanced multimode beam splitter.
+[74] discusses a multi-photon extension of Ref.
+[2].
+This is complementary to Ref.
+[75], which extended the analysis to distributing more ground photons, where parallel-distributed ground photons interfere with the astronomical photon in a balanced multimode beam splitter.
 Ref.
@@
 Consider the simplest unary encoding where there is one memory qubit for each time bin.
-Given $M$ bins, after a single photon is stored, in the 5th time bin, for example, the memories will be in the state
+Suppose there are $M$ time bins in total.
+After a single photon is stored, in the 5th time bin, for example, the memories will be in the state
 $\frac{1}{\sqrt{2}}(\ket{000010...}_{A}\ket{000000...}+$
@@
 VI.4 Unary encoding with binary search
-Figure 6: Schematic for binary search for the time bin into which the photon arrived. This protocol is a new result.
-We can use an encoding that is efficient in the entanglement consumption, where the encoding is akin to a binary search. We choose a block size $M$ where $M\epsilon\sim 1$.
-  • For each run of the parity check, we divide the memory block into two (Fig. 6).
+Figure 6: Schematic for binary search for the time bin into which the photon arrived.
+We can use an encoding that is efficient in the entanglement consumption, where the encoding is akin to a binary search.
+We can choose a block size $M$ such that $M\epsilon\sim 1$.
+For each run of the parity check, we divide the memory block into two (Fig.
+6).
 To search for where the photon is within the time bin, we apply CZ gates between all of Alice’s or Bob’s memory qubits.
-In our example with 4 time bins, a photon arrives in the second bin and is shared non-locally between Alice and Bob. The state is
-$\rho\propto$ |$\ket{0,0}_{A_{1}B_{1}}(\ket{01}_{A_{2}B_{2}}\pm e^{i\phi}\ket{10}_{A_{2}B_{2}})$
-|$\ket{0,0}_{A_{3}B_{3}}\ket{0,0}_{A_{4}B_{4}}$ (33)
-We divide the block of 4 into two halves. The first half is
-$\rho_{1,2}\propto$ |$\ket{0,0}_{A_{1}B_{1}}(\ket{01}_{A_{2}B_{2}}\pm e^{i\phi}\ket{10}_{A_{2}B_{2}})$ (34)
+In our example with 4 bins, a photon arrives in the second bin and is shared non-locally between Alice and Bob. The state is
+$\rho=$ |$\ket{0,0}\bra{0,0}_{A_{1}B_{1}}\otimes\rho^{\star^{\prime}}_{A_{2}B_{2}}\otimes$
+[…]
+$\rho^{\star^{\prime}}$ $=\left(\frac{1+\gamma}{2}\right)\ket{\psi_{+}^{\phi}}\bra{\psi_{+}^{\phi}}+\left(\frac{1-\gamma}{2}\right)\ket{\psi_{-}^{\phi}}\bra{\psi_{-}^{\phi}}$ | (32)
+We divide the block of 4 time bins into two halves. The first half is
+$\rho_{1,2}=$ |$\ket{0,0}\bra{0,0}_{A_{1}B_{1}}\otimes\rho^{\star^{\prime}}_{A_{2}B_{2}}$ (33)
 To perform the check, for all qubits in the half-block, CZ gates are applied between the memory qubit and the shared Bell pair.
-$\rho_{1,2}\otimes\ket{\Phi^{+}}$ $\stackrel{{\scriptstyle 4\times\text{CZ}}}{{\rightarrow}}\rho_{1,2}\times\ket{\Phi^{-}}$ (35)
+$\rho_{1,2}\otimes\ket{\Phi^{+}}\bra{\Phi^{+}}$ | $\stackrel{{\scriptstyle 4\times\text{CZ}}}{{\rightarrow}}\rho_{1,2}\otimes\ket{\Phi^{-}}\bra{\Phi^{-}}$ | (34)
 If there is no photon in the block, such as bins 3 and 4, then the Bell state remains the same.
-$\rho_{3,4}\otimes\ket{\Phi^{+}}$ $\stackrel{{\scriptstyle 4\times\text{CZ}}}{{\rightarrow}}\rho_{3,4}\times\ket{\Phi^{+}}$ (36)
+$\rho_{3,4}\otimes\ket{\Phi^{+}}\bra{\Phi^{+}}$ | $\stackrel{{\scriptstyle 4\times\text{CZ}}}{{\rightarrow}}\rho_{3,4}\otimes\ket{\Phi^{+}}\bra{\Phi^{+}}$ | (35)
 The procedure continues by iteratively dividing the block containing the photon into halves, until a single time bin remains.
-This encoding is still linear in the memory requirements, where the number of memory qubits scales as $1/\epsilon$.
-However, the entanglement consumption has been reduced to scaling with $\sim~{}\log_{2}(1/\epsilon)$
+This encoding is linear in the memory requirements, where the number of memory qubits scales as $1/\epsilon$.
+However, the entanglement consumption has been reduced to $\sim\log_{2}(1/\epsilon)$.
 VI.5 Binary encoding: optical interferometry with quantum Networks [3]
-The binary encoding proposed in [3] is efficient in the entanglement consumption in the parity check. Let us
+The binary encoding proposed in [3] is efficient in memory and entanglement consumption. In this protocol, we
   • assume approximately one photon arrives over $M$ time bins, which $M\sim 1/\epsilon$.
@@
     |$\ket{\bar{1}_{5}}\equiv\ket{1010...0}$
-A encoded logical CNOT is performed between the photonic degrees of freedom and the memory qubits at each time bin
+A encoded logical CNOT (CXm) is performed between the photonic degrees of freedom and the appropriate memory qubits at each time bin
 |$\ket{0}(\ket{\bar{0}},\ket{\bar{1}_{j}})$ $\stackrel{{\scriptstyle\bar{\text{CX}_{m}}}}{{\rightarrow}}\ket{0}(\ket{\bar{0}},\ket{\bar{1}_{j}})$ (36)
@@
 This encoding keeps track of the time-of-arrival information of the photon: empty time bins leave the memories unchanged.
-$\bar{\rho}_{0}\equiv\ket{\bar{0},\bar{0}}\bra{\bar{0},\bar{0}}$ | (39)
-Performing the CX operation,
+Define $\bar{\rho}_{0}\equiv\ket{\bar{0},\bar{0}}\bra{\bar{0},\bar{0}}$ | Perform the CXm operation, we have
 $\rho_{AB}\otimes\bar{\rho}_{0}$ $\stackrel{{\scriptstyle\bar{\text{CX}_{m}}}}{{\rightarrow}}(1-\epsilon)\rho_{\text{vac}}\otimes\bar{\rho}_{0}$
@@
 Then the photonic degree of freedom is decoupled from the memory qubit by measuring it in the X basis.
-(This step is likely challenging, which is solved in the next paper).
-Then, applying the appropriate correction (or post-processing), the memory qubits end up being
-$\rho_{\star}=(1-\epsilon)\bar{\rho}_{0}+$ $\frac{\epsilon(1+|g|)}{2}\ket{\psi^{+}_{\phi,m}}\bra{\psi^{+}_{\phi,m}}$ |
+This step is likely challenging, which can be solved in the next protocol.
+Then, applying the appropriate correction, the memory qubits end up in the state
+$\rho_{AB}\rightarrow(1-\epsilon)\bar{\rho}_{0}+$ $\frac{\epsilon(1+|g|)}{2}\ket{\psi^{+}_{\phi,m}}\bra{\psi^{+}_{\phi,m}}$ |
 $+$ $\frac{\epsilon(1-|g|)}{2}\ket{\psi^{-}_{\phi,m}}\bra{\psi^{-}_{\phi,m}}$ | (40)
-where |$\ket{\psi^{\pm}_{\phi,m}}=\frac{1}{\sqrt{2}}(\ket{\bar{0},\bar{1}_{m}}\pm e^{i\phi}\ket{\bar{1}_{m},\bar{0}})$
-In summary, if a photon arrives in the $(m-1)$-th time bin, then it is stored in such a way that the time bin is flipped according to |$\ket{\bar{1}_{m}}$.
-This protocol uses the same set of memory for each time bin, which means that gates are applied to configure each subsequent time bin encoding.
-Given the short coherence time for each time bin, the gate speed needs to catch up with this operation.
-VI.5.1 Error due to multiphoton contribution
-Let us examine errors due to multiphoton contribution and how to alleviate them.
-Starlight is thermal, and the block size is chosen such that $M\epsilon\sim 1$.
-If the average photon number in the block is 1, then the probability of having more than 1 photon in the block is also high.
-Let us examine the state post-encoding; using the notation of Ref.
-[3]
+where |$\ket{\psi^{\pm}_{\phi,m}}=\frac{1}{\sqrt{2}}(\ket{\bar{0},\bar{1}_{m}}\pm e^{i\phi}\ket{\bar{1}_{m},\bar{0}})$.
+In summary, if a photon arrives in the $(m-1)$-th time bin, then it is stored in such a way that the time bin is prepared according to |$\ket{\bar{1}_{m}}$.
+This protocol uses the same set of memory qubits for all time bins, which require gates to be applied in between each subsequent time bin encoding: the implication is that the gate speed needs to catch up with the time bin size.
+Errors due to multiphoton contribution
+Let us examine errors due to multiphoton contribution in this protocol and how to alleviate them.
+Starlight is thermal, if we choose the time bin block size such that $M\epsilon\sim 1$, then the mean photon number in the block is 1; if so, the probability of having more than 1 photon in the block is also non-negligible.
+Examine the state post-encoding:
 $\rho_{AB}=p_{0}\rho_{\text{vac}}+p_{1}\rho_{1}+(1-p_{0}-p_{1})\rho_{\text{multi}}$ (41)
-where $p_{0}$ is the probability of receiving the vacuum, $p_{1}$ is the probability of having a single photon; and $\rho_{\text{multi}}$ is the state of the memory where more than one photon has landed.
+where $p_{0}$ is the probability of receiving the vacuum, $p_{1}$ is the probability of having a single photon, and $\rho_{\text{multi}}$ is the state of the memory where more than one photon has landed.
 In the case of $\rho_{\text{multi}}$, the memory qubit is considered completely depolarised.
-For $\gamma=1$, we have [3],
-$p_{0}=\frac{1}{\epsilon+1},\qquad p_{1}=\frac{\epsilon}{(1+\epsilon)^{2}}$ (44)
+For $\gamma=1$, $p_{0}={1}/(\epsilon+1)$, $p_{1}={\epsilon}/{(1+\epsilon)^{2}}$ [3].
 If we have $\rho_{AB}^{\otimes M}$ in a block of $M$, then final decoded state in the readout qubit has coefficients given by a trinomial distribution of the input coefficients:
@@
 $\rho=c\rho_{1}+(1-c)\rho_{\text{mix}},\qquad c=\frac{M\epsilon}{[(1+\epsilon)^{M}-1](1+\epsilon)}$ (43)
-This presents a problem, because for high-contrast $c$ needs to be as close to 1 as possible, and this is fundamental; $c$ depends on how we choose $M\epsilon$ (Fig. 7).
+This presents a problem, because for high-contrast $c$ needs to be as close to 1 as possible, and this is fundamental; $c$ depends on how we choose $M\epsilon$ (Fig.
+7).
+If we frequency multiplex the signal such that it is spread over more memory qubits, the probability of infidelity decreases.
+Entanglement consumption rate
 For this protocol, the number of required memory qubits is approximately $\log_{2}(1/\epsilon)$ [4]. Let us calculate the distributed entanglement (|$\ket{\Phi}$) consumption rate:
@@
 $\lambda$ $\Delta\nu$ (Hz) $\Delta\lambda$ $\epsilon$ $\Delta\nu\epsilon\log_{2}(1/\epsilon)$ per sec
-555 nm 10 GHz $10^{-11}$m $10^{-7}$ $23\times 10^{3}$
-555 nm 10 GHz $10^{-11}$m $10^{-10}$ 33
 555 nm 1 THz 1 nm $10^{-7}$ $2\times 10^{6}$
@@
 $\rho\approx(1-\epsilon)\rho_{\text{vac}}+\frac{\epsilon}{2}(\rho_{\nu 1}+\rho_{\nu 2})$ (44)
-$\rho_{\nu 1}=$ $|\displaystyle\frac{1}{2}(\ket{0,1}\bra{0,1}+\ket{1,0}\bra{1,0}+g_{1}\ket{0,1}\bra{1,0}$ |
-$+$ $g_{1}^{*}\ket{1,0}\bra{0,1})_{A,B,\nu 1}\otimes\ket{0,0}\bra{0,0}_{A,B,\nu 2}$
-$\rho_{\nu 2}=$ |$\ket{0,0}\bra{0,0}_{A,B,\nu 1}\otimes\frac{1}{2}(\ket{0,1}\bra{0,1}+$
-[…]
-The notation here is:
-|$\ket{\underbrace{101}_{\text{time}}\underbrace{1}_{\text{freq}},0000}$ (49)
-The first 3 memory qubits encode the time bin, and the 4th encodes the frequency band.
-In this particular example, integrating the stellar photon into the memory leads to the following transformation.
-|$\ket{0,0}_{\nu 1}\ket{0,0}_{\nu 2}\ket{0000_{A},0000_{B}}$
-$\rightarrow$ |$\ket{0,0}_{\nu 1}\ket{0,0}_{\nu 2}\ket{0000_{A},0000_{B}},$ (50)
-|$\ket{1,0}_{\nu 1}\ket{0,0}_{\nu 2}\ket{0000_{A},0000_{B}}$
-$\rightarrow$ |$\ket{1,0}_{\nu 1}\ket{0,0}_{\nu 2}\ket{1010_{A},0000_{B}},$ (51)
-|$\ket{0,1}_{\nu 1}\ket{0,0}_{\nu 2}\ket{0000_{A},0000_{B}}$
-$\rightarrow$ |$\ket{0,0}_{\nu 1}\ket{0,1}_{\nu 2}\ket{0000_{A},1010_{B}},$ (52)
-|$\ket{0,0}_{\nu 1}\ket{1,0}_{\nu 2}\ket{0000_{A},0000_{B}}$
-$\rightarrow$ |$\ket{0,0}_{\nu 1}\ket{1,0}_{\nu 2}\ket{1011_{A},0000_{B}},$ (53)
-|$\ket{0,0}_{\nu 1}\ket{0,1}_{\nu 2}\ket{0000_{A},0000_{B}}$
-$\rightarrow$ |$\ket{0,0}_{\nu 1}\ket{0,1}_{\nu 2}\ket{0000_{A},1011_{B}}.$ (54)
+The notation here for the 2nd frequency band is:
+|$\underbrace{\ket{0_{A_{5}},0_{B_{5}}}}_{\text{freq 1}}\underbrace{\ket{1_{A_{5}},0_{B_{5}}}}_{\text{freq 2}}\underbrace{\ket{1011_{A},0000_{B}}}_{\text{memories}}$
+|$\underbrace{\ket{0_{A_{5}},0_{B_{5}}}}_{\text{freq 1}}\underbrace{\ket{0_{A_{5}},1_{B_{5}}}}_{\text{freq 2}}\underbrace{\ket{0000_{A},1011_{B}}}_{\text{memories}}.$ (45)
+The first 3 memory qubits encode the time bin, and the 4th encodes the frequency band:
+|$\ket{\underbrace{101}_{\text{time}}\underbrace{1}_{\text{freq}}{}_{A},0000_{B}}$ (46)
 This variation has parallel operation over frequencies, at the expense of memory scaling as $R\log_{2}M$, but the entanglement consumption scales logarithmically with $R$ and $M$.
-One technical difficulty associated with using the time bin encoding is that we need to encode the memory qubits as fast as the time bins.
-This timescale is extremely fast compared to the current gate times for most platforms.
 VI.7 Incorporating quantum error correction [5]
@@
 Note that the excited state |$\ket{e}$ is not populated.
-The technically difficult part of this protocol includes:
-  • The STIRAP step needs to be extremely low noise: the stellar state is unencoded at this point, and the error rate must be $\ll\epsilon$.
-  • For a complete population transfer, here $t=50$, which appears long. The cavity quality needs to be sufficiently high to retain the stellar photon until the transfer is complete.
-  • STIRAP schemes have very narrow bandwidth. (Rare earth: 10k Hz)
+The technically difficult part of this protocol may include: 1) The STIRAP step needs to be extremely low noise: the stellar state is unencoded at this point, and the error rate must be $\ll\epsilon$.
+2) For a complete population transfer, here $t=50$, which appears long.
+The cavity quality needs to be sufficiently high to retain the stellar photon until the transfer is complete.
+e) STIRAP schemes tend to have very narrow bandwidth, in rare Earth ions the typical bandwidth is around 10k Hz
 As an example of a suitable STIRAP scheme, consider a 87Rb atom trapped in an optical tweezer and coupled to a fibre Fabry-Perot cavity as recently demonstrated [78].
@@
 We assume a cavity finesse $F\approx 2\times 10^{5}$, a waist radius ${\rm w}_{r}\approx 2\,\mu$m and a length $L\approx 40\,\mu$m resulting in a cooperativity of $C=3\lambda^{2}F/(2\pi^{3}{\rm w}_{r}^{2})\approx 1500$ (as demonstrated in [80]) with a coupling strength of $g=\sqrt{3\lambda^{2}C\gamma}/(2\pi^{2}{\rm w}_{r}^{2}L)\approx 2\pi\times 400\,$MHz and $\kappa=\pi C/LF\approx 2\pi\times 20\,$MHz (FWHM), so that $\gamma/\kappa\approx 0.3$.
-Over a population transfer period of $T$ one estimates a fidelity of $F=e^{-T\kappa}$ which for $T=50g^{-1}$ would be too small, $F<0.1$.
-Better cavities providing a factor of $10$ improvement in $g/\kappa$ would provide $F>0.75$.
-Note Tweezer induced dephasing rates on state |$\ket{1}$ can be treated as negligible $\gamma_{\phi}/g=0.03\times 10^{-6}$ [77].
+Tweezer-induced dephasing rates on state |$\ket{1}$ can be treated as negligible $\gamma_{\phi}/g=0.03\times 10^{-6}$ [81], so the dominant source of error would be cavity decay.
+Over a population transfer period of $T$, with on average $1/2$ photon occupation in the cavity, one estimates a fidelity of $F=e^{-T\kappa/2}$ which for $T=50g^{-1}$ would imply $F\approx 0.29$.
+Better cavities providing a factor of $5$ improvement in $g/\kappa$ would enable $F>0.77$, which could be further improved by post-selecting on null results from single photon detectors monitoring leaked photons [82].
 VI.8 Continuous-variable investigations
@@
 Practical difficulties further impede achieving any quantum advantage, limiting the merits of a CV teleportation-based strategy for stellar interferometry.
-Conclusion: TMSV teleportation-based strategies are practically infeasible.
+Consistent with Ref.
+[69], we conclude that TMSV teleportation-based strategies are practically infeasible if there is no quantum repeater available.
 VI.9 Vacccum beam guide for quantum networks [6]
@@
 The system avoids the complexity of quantum repeaters by requiring only passive optical elements.
-Key sources of loss—lens imperfections, residual gas absorption, and beam misalignment—are rigorously analysed and shown to be manageable with current technology.
-The system avoids the complexity of quantum repeaters by requiring only passive optical elements.
-Key sources of loss—lens imperfections, residual gas absorption, and beam misalignment—are rigorously analysed and shown to be manageable with current technology.
-This architecture is not only promising for quantum communication but also enables transformative advances in quantum-enabled long-baseline optical interferometry (QVLBI).
-In particular, we see two distinct, yet complementary, directions in which the VBG infrastructure could be adapted for interferometric science.
-VI.9.1 High-Efficiency Entanglement Distribution for Quantum Interferometry
-In a repeater-based VLBI architecture, VBGs could serve as the physical backbone for entanglement distribution between telescope nodes. This offers several key advantages:
+Key sources of loss—lens imperfections, residual air absorption, and beam misalignment—are analysed and shown to be manageable with current technology.
+This architecture is not only promising for quantum communication but also enables advances in quantum-enabled O-VLBI.
+In particular, we see two distinct, yet complementary, directions in which the VBG infrastructure could be adapted:
+VI.9.1 High-Efficiency Entanglement Distributor
+In a repeater-based VLBI architecture, VBGs could serve as the physical backbone for entanglement distribution between telescope nodes.
+This offers several advantages.
 Low-loss transmission over thousands of kilometres enables high-rate distribution of entangled photons with reduced need for quantum repeaters.
@@
 VI.9.2 Using the VBG as an interferometer arm
-Alternatively, the VBG itself can be directly repurposed as an interferometric guide channel, effectively acting as one arm of a giant optical interferometer:
+Alternatively, the VBG itself can be directly repurposed as an interferometric guide channel, effectively acting as one arm of a large optical interferometer.
 In this approach, starlight collected at separate telescopes is coupled into the VBG and transported to a common detection point for interference-based measurement.
@@
 This could be achieved through delay lines and phase-locking mechanisms, possibly guided by a bright reference source or classical beacon.
-This direct-interference approach would closely mimic classical optical interferometry but at scales (1000+ km) previously accessible only in the radio domain.
 In addition, if implemented on the Moon, the vacuum beam guide would benefit from the natural hard vacuum of the lunar environment, eliminating the need for complex evacuated tubes.
-This could significantly reduce engineering overhead while enabling ultra-long baseline quantum interferometry with minimal optical loss.
+This could significantly reduce engineering overhead with minimal optical loss.
 VII Open questions
@@
 STIRAP) and faster dynamical techniques?
-Adiabatic loading schemes (e.g.
-STIRAP) offer robustness but are slow, while dynamical schemes allow faster coupling at the cost of greater sensitivity to timing and control.
+Adiabatic loading schemes offer robustness but are slow, while dynamical schemes allow faster coupling at the cost of having high noise to timing and control.
 The optimal choice depends on memory coherence time and system timing constraints.
@@
 For binary encoding, the number of memory qubits scales as $\log(1/\epsilon)$, but practical constraints such as detector dead time and memory reset times also play a role.
-Is temporal or wavelength multiplexing unavoidable?
-If so, how can we implement it without overwhelming the memory and feedforward control?
-Can we design architectures that minimise multiplexing while preserving sensitivity?
-While multiplexing boosts photon capture and entanglement rates, it increases system complexity.
-Minimising multiplexing while maintaining detection efficiency remains an open design question.
-In summary, mode-resolved quantum VLBI imposes a unique set of challenges that bridge quantum optics, temporal mode engineering, and real-time feedback.
-Solving these problems will be essential for unlocking the full potential of quantum-enhanced astronomical imaging and precision geodesy.
+How can we most efficiently implement multiplexing without overwhelming the memory and feedforward control?
+The atmosphere fluctuates on the timescale of milliseconds.
+What is the best way of correcting for the phase variations?
+What requirements are needed for a potential artificial guiding star?
+Can we collect enough photons by using broadband light to avoid using a guiding star?
+How might established techniques such as closure-phase observables be adapted within quantum-enhanced architectures?
+More broadly, can quantum estimation theory enable new imaging schemes that go beyond pairwise visibility measurements and the standard van Cittert–Zernike framework?
+In the absence of full quantum repeaters, what level of VLBI enhancement is realistically achievable using repeater-less quantum links?
+Can the distribution of additional ground photons, particularly with more robust temporal or mode entanglement, meaningfully improve signal-to-noise ratio?
+Furthermore, how might such protocols connect with or inform developments in all-optical quantum repeater architectures?
+Beyond classical Fisher information and maximum likelihood methods, could quantum-enhanced or learning-assisted inference techniques play a role in astronomical imaging?
+Recent work at the interface of quantum estimation and learning theory?
+E.g., Ref.
+[86] suggests a potentially fruitful direction for developing robust, data-efficient reconstruction techniques tailored to interferometric observations.
 VIII Concluding remarks
-This report has outlined the feasibility and potential of quantum-enabled optical Very Long Baseline Interferometry (VLBI).
-By integrating quantum information protocols, many long-standing challenges in optical VLBI can be mitigated.
-From a technological perspective, we examined a range of architectures, including repeater-based entanglement distribution, binary-encoded quantum memories, and hybrid protocols.
-We believe the uses of the vacuum beam guide architecture, either as a lossless medium for entanglement distribution or as a coherent optical delay channel for direct interferometry, offer scalable paths toward kilometre-scale optical baselines.
+This report has outlined the feasibility and potential of quantum-enabled optical VLBI. By integrating quantum information protocols, many long-standing challenges in optical VLBI can be mitigated.
+From a technological perspective, we examined a range of architectures.
 Encouragingly, most of the individual hardware components—such as high-cooperativity optical cavities, long-lived quantum memories, and entangled photon sources—are now approaching the threshold of practical feasibility.
@@
 In particular, time-bin multiplexed architectures impose stringent requirements on gate execution times and error rates that exceed current capabilities in most physical platforms.
-These limitations currently constrain the implementation of real-time encoding, decoding, and error correction procedures necessary for robust operation.
-Nevertheless, the convergence of quantum technologies toward these benchmarks is rapid.
+Nevertheless, the development of quantum technologies toward these benchmarks is rapid.
 Continued improvements in quantum control and photonic integration may soon enable the deployment of quantum-enhanced VLBI systems capable of achieving diffraction-limited resolution at optical wavelengths.
@@
 Acknowledgements.
-We thank Max Charles, John Bartholomew, Michael Steel, Christian Schwab, John Jeffers, Lorenzo Maccone, and Chiara Macchiavello for insightful discussions.
+We thank Max Charles, John Bartholomew, Michael Steel, Christian Schwab, John Jeffers, Lorenzo Maccone, and Chiara Macchiavello, Yujie Zhang, Yunkai Wang and Virginia Lorenz for insightful discussions.
 ZH is supported by an ARC DECRA Fellowship (DE230100144) “Quantum-enabled super-resolution imaging” and the EPSRC International Network in Space Quantum Technologies (EP/W027011/1).
@@
 We pay our respects to Elders past, present and future.
-Appendix A Tables for reference
-Wavelength $\lambda$ (nm) Frequency $\nu$ (THz) $\Delta\lambda$ (pm)
-600 500 0.12
-1500 200 0.75
-10000 30 33
-Table 4: Conversion of a frequency bandwidth of $\Delta\nu=10^{8}$ Hz to wavelength bandwidth at different central wavelengths. The coherence time is $10$ ns, as per Eq. (19).
-Wavelength $\lambda$ (nm) Frequency $\nu$ (THz) $\Delta\nu$ (THz)
-400 750 18.8
-500 600 12.0
-600 500 8.3
-1000 300 3.0
-1550 193.5 1.25
-2000 150 0.75
-Table 5: Conversion of a 10 nm wavelength bandwidth to frequency bandwidth at different central wavelengths.
-Appendix B Calculating the mean photon numbers of nearby stars and exoplanets
+Appendix A Calculating the mean photon numbers of nearby stars and exoplanets
 Planet mass ($M_{\oplus}$) temperature (K) distance (light year) host apparant brightness ($m_{AB}$)
@@
 In this section, we include details of how we arrive at the mean photon numbers of nearby stars and exoplanets.
-Here the monochromatic absolute magnitude $m_{AB}$ is defined as the logarithm of a spectral flux density with the usual scaling of astronomical magnitudes The zero-point is approximately about 3631 Jy, where
+Here the monochromatic absolute magnitude $m_{AB}$ is defined as the logarithm of a spectral flux density with the usual scaling of astronomical magnitudes.
+The zero-point is approximately about 3631 Jy, where
 […]
 […]
-Denote $F_{\nu}$ as the spectral flux density
+Denote $F_{\nu}$ as the spectral flux density.
 $m_{\text{AB}}$ $\approx-2.5\log_{10}\left(\frac{F_{\nu}}{3631~{}\text{Jy}}\right)$ (51)
@@
 $\Phi_{\lambda}=\frac{F_{\lambda}}{hc/\lambda}=\frac{F_{\nu}}{h\times\lambda}$ (57)
-Now, to calculate the total photon specific bandwidth, we assume the density is approximately constant over the bandwidth (this holds for narrow bandwidths):
+Now, to calculate the total photon-specific bandwidth, we assume the density is approximately constant over the bandwidth (this holds for narrow bandwidths):
 $\Phi_{\text{tot}}$ $=\Phi_{\lambda}\times\Delta\lambda$ (58)
@@
 […]
-For the CHARA array, visible interferometry has bandwith between 10-50 nm.
+For the CHARA array, visible interferometry has a bandwidth between 10-50 nm.
 Putting in the numbers, $\Delta\nu=10\times 10^{9}$, $\lambda=555\times 10^{-9},m=10$
@@
 13 1.65 $\mu$m 1 nm 100 GHz 2095 $10^{-12}$ $1.9\times 10^{-8}$ $1.9\times 10^{-17}$
-Table 7: Expected number of photons to collect at given wavelength, bandwidth for the host star.
+Table 5: Expected number of photons to collect at a given wavelength, bandwidth for the host star.
 The wavelengths of interest are: 760nm is the absorption line for molecular oxygen; for Earth-like conditions the pressure-broadening linewidth is typically $\sim$ 2 pm at FWHM; 1.65 $\mu$m is a methane absorption line, and for earth-like conditions the FWHM ×$\times 10^{-4}\mu$m.
 Here we have assumed a collection area of 10 m2.
-Spectral demultiplexing doesn’t decrease $\epsilon$ due to the fact that that the effective time bins get larger.
-We assume a conservative $10^{-9}$ factor between the star and planet’s brightness.
+Spectral demultiplexing doesn’t decrease $\epsilon$ due to the fact that the effective time bins get larger.
+We assume a conservative $10^{-9}$ factor between the star and the planet’s brightness.
 These numbers might differ depending on what assumption is made about the temperature of the star.